### v2

# DMSC: Dynamic Multi-Scale Coordination Framework for Time Series Forecasting

Abstract
Time Series Forecasting (TSF) faces persistent challenges in modeling intricate temporal dependencies across different scales. Despite recent advances leveraging different decomposition operations and novel architectures based on CNN, MLP or Transformer, existing methods still struggle with static decomposition strategies, fragmented dependency modeling, and inflexible fusion mechanisms, limiting their ability to model intricate temporal dependencies. To explicitly solve the mentioned three problems respectively, we propose a novel Dynamic Multi-Scale Coordination Framework (DMSC) with Multi-Scale Patch Decomposition block (EMPD), Triad Interaction Block (TIB) and Adaptive Scale Routing MoE block (ASR-MoE). Specifically, EMPD is designed as a built-in component to dynamically segment sequences into hierarchical patches with exponentially scaled granularities, eliminating predefined scale constraints through input-adaptive patch adjustment. TIB then jointly models intra-patch, inter-patch, and cross-variable dependencies within each layer’s decomposed representations. EMPD and TIB are jointly integrated into layers forming a multi-layer progressive cascade architecture, where coarse-grained representations from earlier layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways. And ASR-MoE dynamically fuses multi-scale predictions by leveraging specialized global and local experts with temporal-aware weighting. Comprehensive experiments on thirteen real-world benchmarks demonstrate that DMSC consistently maintains state-of-the-art (SOTA) performance and superior computational efficiency for TSF tasks. Code is available at https://github.com/1327679995/DMSC.
Introduction
Time Series Forecasting (TSF) constitutes a pivotal capability across numerous fields such as energy consumption(Alvarez et al. 2010; Guo et al. 2015), healthcare monitoring(Tran, Nguyen, and Shahabi 2019; Wei et al. 2022), transportation scheduling(Guo et al. 2020; Jin et al. 2021), weather forecasting(Wu et al. 2023b; Bi et al. 2023), economics(Chen et al. 2023b; Yu et al. 2023) and urban computing(Li et al. 2024; Yuan et al. 2024). Time series data exhibit intricate temporal dependencies, encompassing nonlinear relationships, long-term and short-term patterns such as trends, seasonality and dynamic variable couplings(Qiu et al. 2024). Owing to non-stationarity(Nason 2006) and complex interconnections(Dagum and Bianconcini 2016), traditional approaches struggle to effectively deal with time series data, making effectively dependency modeling a critical challenge for enhancing predictive precision(Shao et al. 2025). In recent years, the powerful ability of deep learning to capture complex dependency relationships in time series data has been demonstrated, making it a formidable tool in the field of TSF(Chen et al. 2023a).
However, intricate dependencies in time series inherently manifest across multiscales: coarser granularities encapsulate long-term trends while finer resolutions capture short-term fluctuations and periodic patterns. Existing single-scale modeling approaches often fail to balance fine-grained local details and holistic global trends(Wang et al. 2025). To effectively model multiscale temporal dependencies, recent research has pioneered multiscale modeling frameworks that concurrently extract global and local patterns, thereby mitigating the representational limitations of single-scale approaches. While previous implementations often employ moving average downsampling(Wang et al. 2024b; Huang et al. 2025), which often remains insufficient in capturing local semantic information. In contrast, more effective alternatives focus on patch-wise and variable-wise decomposition operations that decompose sequence across feature and temporal dimensions. Such decomposition operations better preserves both global trends and local semantics, and multi-length patches inherently learn multi-scale representations. Critically, prevailing methods exhibit two constraints:1) dependence on fixed-scale decomposition strategies that lack dynamic scale adaptation; 2) fragmented processing of temporal dependencies and cross-variable interactions, undermining holistic dependency modeling.
The core value of multiscale feature extraction resides in mining complementary information across diverse temporal granularities and hierarchical levels, thereby overcoming the limitations of single-scale modeling. Following feature capture, the effective design of prediction heads for multiscale fusion is equally important. Conventional approaches typically employ either linear projection layers(Li, Li, and Diao 2025), additive combinations of multiple projections(Wang et al. 2024b), or spectral amplitude-weighted aggregation(Wu et al. 2023a). However, above simplistic fusion method exhibit fundamental inadequacies: they fail to leverage heterogeneous dominance patterns across multiscales, disregard dynamic inter-scale importance, and induce quadratic parameter growth with increasing input lengths and scale counts, severely compromising deployment efficiency(Qiu et al. 2024).
To effectively modeling complex temporal dependencies which inherently manifest across multiple patterns and scales, we propose a Dynamic Multi-Scale Coordination framework (DMSC) for TSF. This framework dynamically decomposes time series and extracts features across scales. It effectively captures intricate dependencies across temporal resolutions and cross-variable interactions. Meanwhile, it can adaptively predict learned representations through temporal-aware fusion. Specifically, DMSC processes intricate dependencies through a multi-layer progressive architecture and Adaptive Scale Routing Mixture-of-Experts (ASR-MoE). The multi-layer progressive cascade architecture incorporates Embedded Multi-Scale Patch Decomposition Block (EMPD), Triad Interaction Block (TIB), where EMPD transforms input into 3D representations at varying granularities and TIB models heterogeneous dependencies within each layer. In this architecture, coarse-grained representations from shallow layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways. After multi-scale feature extraction, ASR-MoE is employed to perform time-aware weighted dynamic fusion for multi-scale prediction, enabling adaptive integration of features across scales. By synergistically integrating dynamic patch decomposition, deeper models and sparse MoE principles, DMSC achieves full-spectrum multi-scale coordination across embedding, feature extraction, and prediction stages. This dynamic co-optimization enable it to achieve SOTA performance across 13 real-world benchmarks while maintaining high efficiency and low computational complexity. Our contributions are as follows:
  • A novel multi-layer progressive cascade architecture is designed to jointly integrate Embedded Multi-Scale Patch Decomposition Block (EMPD) and Triad Interaction Block (TIB) across layers. Unlike fixed-scale approaches, EMPD dynamically adjusts patch granularities by employing a lightweight neural controller based on input characteristics, and TIB jointly models intra-patch, inter-patch, and inter-variable dependencies through gated feature fusion to form a coarse-to-fine feature pyramid, where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated residual pathways.
  • Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) is proposed to resolve static fusion limitations in prediction. ASR-MoE establishes a hierarchical expert architecture to explicitly decouples long-term and short-term dependencies, where global-shared experts capture common long-term temporal dependencies, while local-specialized experts model different short-term variations. Significantly, a temporal-aware weighting aggregator is designed to dynamically compute scale-specific prediction contributions with historical memory.
  • Extensive experiments demonstrate that DMSC achieves SOTA performance and superior efficiency across multiple TSF benchmarks.
Related Work
Time Series Forecasting Models
Existing deep models for TSF tasks can be broadly categorized into MLP-based, CNN-based, and Transformer-based architectures(Wang et al. 2024c). MLP-based models(Oreshkin et al. 2020; Yi et al. 2023; Chi et al. 2021) primarily model temporal dependencies through predefined decomposition schemes and fully-connected architectures. While demonstrating simplicity and efficiency in capturing intra-series patterns, they exhibit critical limitations in modeling long-term dependencies and complex inter-series relationships. In contrast, CNN-based models(Bai, Kolter, and Koltun 2018; LIU et al. 2022; Wang et al. 2023) leverage different convolutions to effectively capture local dependencies. Although they remain constrained in modeling global relationships, recent efforts have mitigated this through large-kernel convolutions(Luo and Wang 2024) and frequency-domain analysis(Wu et al. 2023a). Nevertheless, Transformer-based models(Zhou et al. 2021; Wu et al. 2021; Wang et al. 2024a) have emerged as a powerful paradigm by utilizing self-attention mechanisms(Vaswani et al. 2017) to model superior long-term dependency across long sequences, so that effectively capturing persistent temporal relationships. However, Transformer-based models face substantive criticism regarding permutation invariance and quadratic computational complexity limitations in recent researches(Zeng et al. 2023; Tang and Zhang 2024). While significant progress has been made, existing deep architectures primarily operate on single or fixed decomposition, and often model dependencies in a fragmented manner, lacking a mechanism for dynamic, adaptive multi-scale decomposition and coordinated dependency modeling.
Decomposition Operations in TSF
To effectively address intricate temporal dependencies, most studies have adopted multiscale modeling perspective, employing diverse multiscale decomposition operations on time series. TimesNet(Wu et al. 2023a) learns frequency-adaptive periods in the frequency domain, transforming time series into 2D tensors to explicitly model intra-period and inter-period variations. TimeMixer(Wang et al. 2024b) decomposes sequences into seasonal and trend components(Wu et al. 2021), subsequently applying different granularity mixing strategies to integrate multiscale features. However, simply using moving average inadequately preserves local semantic patterns. Diverging from the above operations, some methods focus on variable-wise and patch-wise disentanglement. PatchTST(Nie et al. 2023) segments series into subseries-level patches as input tokens, employing channel independence to capture local semantics. ITransformer(Liu et al. 2024) embeds entire series as singular tokens to model extended global representations. TimeXer(Wang et al. 2024d) hierarchically represents variables through dual disentanglement, variable-wise for cross-channel interactions and patch-wise for endogenous variables. However, these decomposition techniques are constrained by static or predefined scale configurations, which hinders their adaptability to complex temporal patterns.
Figure 1: The framework of DMSC. EMPD dynamically segments input series into hierarchical patches, TIB jointly models three dependencies, and ASR-MoE fuses multiscale predictions adaptively.
Multiscale Prediction Operations in TSF
Meanwhile, researches also focuses on multiscale operations in prediction heads. TimeMixer(Wang et al. 2024b) employs dedicated projectors-one per scale-to generate scale-specific predictions. Diverging from simplistic additive aggregation, TimeMixer++(Wang et al. 2025) utilizes frequency-domain amplitude analysis to weightedly aggregate projector outputs. TimeKAN(Huang et al. 2025) adopts frequency transformation and padding operations to unify multiscale representations into identical dimensions, enabling holistic scale integration. Current multiscale fusion mechanisms predominantly rely on static weighting schemes or fixed designs. They fail to dynamically prioritize different scales with intricate temproal dependencies. In contrast to these fragmented approaches, DMSC pioneers fully-spectrum multiscale coordination spanning embedding, extraction, and prediction stages, achieving optimal balance between computational efficiency and forecasting accuracy.
Dynamic Multi-Scale Coordination Framework
Time series forecasting addresses the fundamental challenge of predicting future values ${Y}={y_{t+1},...,y_{t+\tau}}\in\mathbf{R}^{\tau\times{C}}$ from historical observations ${X}={x_{1},...,x_{t}}\in\mathbf{R}^{t\times{C}}$, where $\tau$ is the prediction horizon, $t$ is the history horizon, and $C$ is the number of variate. To model complex multiscale dependencies inherent in real-world series, the Dynamic Multi-Scale Coordination (DMSC) framework is proposed. As illustrated in Fig. 1, DMSC establishes full-spectrum multi-scale coordination across embedding, feature extraction, and prediction stages using EMPD, TIB and ASR-MoE.
Multi-Layer Progressive Cascade Architecture
The multi-layer progressive cascade architecture enables hierarchical feature learning through stacked EMPD-TIB units, creating a dynamic feature pyramid where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated pathways. Formally, given input $\mathbf{X}\in\mathbf{R}^{C\times L}$, layer $l$ processes:
$\mathcal{T}=\{\mathbf{F}_{1},\mathbf{F}_{2},...,\mathbf{F}_{l}\},$
$\mathbf{F}_{l}=\textbf{TIB}_{l}(\mathbf{Z}_{l}),$
$\mathbf{Z}_{l}=\textbf{EMPD}_{l}(\mathcal{G}_{l}(\mathbf{F}_{l-1})+\mathbf{X}),$ (1)
where $\mathcal{T}$ is the set of hierarchical features generated by this architecture for ASR-MoE, with each $\mathbf{F}_{l}\in\mathbf{R}^{C\times{D}}$ representing the $l$-th layer’s output. $\mathcal{G}_{l}$ is a gated projection matrix that adaptively modulates the residual flow. This cascade flow enables progressive refinement of multiscale representations, where $\mathbf{F}_{l-1}$ adaptively guides $\mathbf{Z}_{l}$ with residual connections, and $\textbf{TIB}_{l}$ iteratively enhances representations by jointly integrating intra-patch, inter-patch, and cross-variable dependencies. By interleaving dynamic patch decomposition (EMPD) with triadic dependency modeling (TIB), our cascade architecture achieves input-aware progressive refinement of temporal granularities, enabling coherent modeling of both short-term dynamics and long-range trends while maintaining parameter efficiency.
Embedded Multi-Scale Patch Decomposition
Patch decomposition serves as a fundamental operation to transform time series into structured representations that preserve local semantics while enabling hierarchical pattern discovery. Unlike conventional single-fixed or fixed-scale approaches, the EMPD block introduces input-adaptive and hierarchical decomposition mechanism as a built-in component, dynamically adjusting patch granularities based on sequence characteristics. Thus eliminating rigid predefined patch configurations while natively integrating multiscale decomposition into the model architecture.
Specifically, given an input time series tensor $\mathbf{X}\in\mathbf{R}^{C\times{L}}$(where $C$ is the number of features, $L$ is the sequence length), EMPD first computes a scale factor $\alpha$ through a lightweight neural controller:
$\alpha=\mathcal{N}_{\theta}(\Phi_{\mathrm{GAP}}(\mathbf{X}))$ (2)
where $\mathcal{N}_{\theta}$ is a micro-MLP with sigmoid activation, $\Phi_{\mathrm{GAP}}$ compresses temporal dimensions through averaging. Factor $\alpha\in[0,1]$ dynamically determines the base patch length, and then the patch length decays exponentially as follows:
$\mathbf{P}_{{base}}=[\mathbf{P}_{min}+\alpha\cdot(\mathbf{P}_{max}-\mathbf{P}_{min})],$
$\mathbf{P}_{l}=\mathrm{max}(\mathbf{P}_{min},[\mathbf{P}_{base}/\tau^{l}]),$ (3)
where $\mathbf{P}_{max}$ and $\mathbf{P}_{min}$ denoting the learned parameter for minimum and maximum patch bounds, and $\mathbf{P}_{l}$ means patch length in layer $l$, which ensures shallow layers process coarse-grained dependencies while deeper layers capture fine-grained patterns. Then EMPD applies replication padding to handle boundary effects, followed by hierarchical patch unfolding. The operations are formalized as:
$\mathbf{X}_{p}^{l}=\mathrm{Unfold}(\mathrm{Padding}_{S_{l}}(X)),$ (4)
where $\mathrm{Padding}_{S_{l}}(\cdot)$ denotes replication padding with stride $S_{l}=P_{l}/2$ and $\mathrm{Unfold}_{P_{l},S_{l}}$ unfolds the padded sequence into patches of length $P_{l}$ with stride $S_{l}$. This operation yields a 3D patch tensor $\mathbf{X}_{u}^{l}\in\mathbf{R}^{C\times{N_{l}}\times{P_{l}}}$. Then EMPD transforms $\mathbf{X}_{u}^{l}$ into a unified embedding space via linear projection:
$\mathbf{Z}^{l}=\mathrm{Projector}(\mathbf{X}_{p}^{l}),$ (5)
This direct linear projection avoids redundant flattening operations, efficiently transforming patch-wise temporal features into a unified embedding space while preserving the hierarchical structure of multiscale patches.
Triad Interaction Block
TIB is designed to model heterogeneous dependencies within the multiscale patch representations generated by EMPD, through joint capture of intra-patch, inter-patch, and cross-variable features. Given the input tensor $\mathbf{Z}^{l}\in{R}^{C\times{N_{l}}\times{D}}$, TIB integrates three complementary dependency types in a coherent framework. Specifically, TIB first processes intra-patch dependencies using depth-wise separable convolutions via $\mathbf{\zeta}_{intra}$ to capture fine-grained local patterns:
$\mathbf{F}_{{intra}}^{l}=\mathbf{\zeta}_{intra}(\mathbf{Z}^{l}),$ (6)
where $\mathbf{\zeta}_{intra}$ consists the depth-wise convolution and $\mathrm{Conv1d}$ to captures local temporal contexts and project features back to the embedding space. This operation preserves intra-patch continuity while extracting localized representations.
For inter-patch dynamics $\mathbf{F}_{{intra}}^{l}$, TIB employs dilated convolutions with adaptive pooling via $\mathbf{\zeta}_{inter}$:
$\mathbf{F}_{{inter}}^{l}=\mathbf{\zeta}_{inter}(\mathbf{F}_{{intra}}^{l}),$ (7)
where $\mathbf{\zeta}_{inter}$ consists the dilated convolution and adaptive pooling to capture broader temporal contexts and patch-level information without increasing computational cost.
Then TIB models dependencies between different features by generating adaptive weighting coefficients $\mathbf{G}^{l}\in\mathbf{R}^{{C}\times{1}}$ from globally aggregated inter-patch features $\bar{\mathbf{F}}_{{inter}}^{l}\in\mathbf{R}^{{1}\times{D}}$ as follows:
$\mathbf{G}^{l}=\sigma\left(\mathrm{MLP}(\bar{\mathbf{F}}_{{inter}}^{l})\right),$
$\mathbf{F}_{{cross}}^{l}=\mathbf{G}^{l}\odot\mathbf{F}_{{inter}}^{l},$ (8)
where $\bar{\mathbf{F}}_{{inter}}^{l}$ denotes the global average of $\mathbf{F}_{{inter}}^{l}$, and the sigmoid function $\sigma$ act as a feature-specific gate. The weights $\mathbf{G}^{l}$ are then applied element-wise to $\mathbf{F}_{{inter}}^{l}$ through multiple operation $\odot$, which adaptively scales each feature’s contribution based on its own global relevance, enabling context-aware cross-variable interactions.
Finally, three dependency representations are adaptively fused via learned gating weights, and the fused features are then integrated with residual connection and normalization:
$\mathbf{F}^{l}_{{fused}}=\sum_{i=1}^{3}\mathbf{G}^{l}_{i}\odot\mathbf{F}^{l}_{i},$
$\mathbf{F}^{l}_{{output}}=\mathrm{LayerNorm}(\mathbf{F}^{l}_{{fused}}+\mathbf{Z}^{l}),$ (9)
where $\mathbf{G}^{l}_{i}$ is learned gating weights for each representations. This architecture enables TIB to dynamically balance contributions of different dependency types based on input characteristics, thus effectively capturing intricate dependencies, providing rich temporal features for multiscale prediction.
Models DMSC(Ours) TimeMixer iTransformer PatchTST Dlinear TimesNet Autoformer TimeXer PatchMLP TimeKAN AMD
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.415 0.426 $0.456$ $0.444$ $0.452$ $0.446$ $0.445$ $0.445$ $0.461$ $0.457$ $0.479$ $0.466$ $0.670$ $0.564$ $0.460$ $0.452$ $0.458$ $0.445$ $\underline{0.429}$ $\underline{0.432}$ $0.452$ $0.439$
ETTh2 0.355 0.383 $\underline{0.372}$ $\underline{0.400}$ $0.384$ $0.407$ $0.383$ $0.412$ $0.563$ $0.519$ $0.411$ $0.421$ $0.488$ $0.494$ $0.374$ $0.404$ $0.406$ $0.422$ $0.389$ $0.409$ $0.616$ $0.558$
ETTm1 0.368 0.383 $0.384$ $0.398$ $0.408$ $0.412$ $0.386$ $0.400$ $0.404$ $0.408$ $0.418$ $0.418$ $0.636$ $0.534$ $0.391$ $0.395$ $0.387$ $0.398$ $\underline{0.376}$ $0.397$ $0.394$ $\underline{0.396}$
ETTm2 0.268 0.317 $0.278$ $0.324$ $0.292$ $0.336$ $0.288$ $0.334$ $0.354$ $0.402$ $0.291$ $0.330$ $0.502$ $0.452$ $\underline{0.277}$ $\underline{0.323}$ $0.287$ $0.329$ $0.282$ $0.330$ $0.288$ $0.332$
Electricity 0.170 0.258 $0.190$ $0.280$ $\underline{0.176}$ $\underline{0.265}$ $0.205$ $0.295$ $0.226$ $0.319$ $0.194$ $0.293$ $0.492$ $0.523$ $0.201$ $0.275$ $0.200$ $0.296$ $0.201$ $0.290$ $0.206$ $0.287$
Exchange 0.336 0.391 $0.359$ $\underline{0.401}$ $0.362$ $0.406$ $0.389$ $0.417$ $\underline{0.339}$ $0.414$ $0.430$ $0.446$ $0.539$ $0.519$ $0.398$ $0.422$ $0.383$ $0.417$ $0.384$ $0.414$ $0.366$ $0.408$
Weather 0.233 0.269 $0.245$ $0.274$ $0.261$ $0.281$ $0.258$ $0.279$ $0.265$ $0.316$ $0.256$ $0.283$ $0.348$ $0.382$ $\underline{0.242}$ $\underline{0.272}$ $0.252$ $0.277$ $0.244$ $0.274$ $0.271$ $0.291$
Traffic 0.407 0.274 $0.509$ $0.307$ $\underline{0.422}$ $\underline{0.283}$ $0.482$ $0.309$ $0.688$ $0.428$ $0.633$ $0.334$ $1.028$ $0.605$ $0.466$ $0.287$ $0.539$ $0.364$ $0.590$ $0.373$ $0.547$ $0.345$
Solar 0.213 0.258 $0.228$ $0.279$ $0.238$ $\underline{0.263}$ $0.249$ $0.292$ $0.330$ $0.401$ $0.268$ $0.286$ $0.727$ $0.634$ $\underline{0.226}$ $0.269$ $0.298$ $0.301$ $0.289$ $0.321$ $0.363$ $0.337$
Table 1: Long-term forecasting results. All the results are averaged from 4 different prediction lengths {96, 192, 336, 720}, and the look-back length is fixed to 96 for all baselines. A lower MSE or MAE indicates a better prediction, with the best in boldface and second in underline.
Models DMSC(Ours) TimeMixer iTransformer PatchTST Dlinear TimesNet Autoformer TimeXer PatchMLP TimeKAN AMD
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03 0.131 0.241 $0.190$ $0.288$ $0.261$ $0.328$ $0.308$ $0.368$ $0.279$ $0.377$ $\underline{0.154}$ $\underline{0.254}$ $0.513$ $0.517$ $0.180$ $0.280$ $0.220$ $0.289$ $0.283$ $0.357$ $0.281$ $0.368$
PEMS04 0.129 0.244 $0.230$ $0.324$ $0.913$ $0.801$ $0.365$ $0.405$ $0.295$ $0.389$ $\underline{0.138}$ $\underline{0.250}$ $0.472$ $0.495$ $0.327$ $0.418$ $0.203$ $0.305$ $0.297$ $0.372$ $0.327$ $0.397$
PEMS07 $\underline{0.092}$ $\underline{0.193}$ $0.165$ $0.256$ $0.109$ $0.214$ $0.438$ $0.422$ $0.329$ $0.394$ $0.109$ $0.216$ $0.418$ $0.466$ 0.088 0.192 $0.166$ $0.262$ $0.264$ $0.346$ $0.196$ $0.318$
PEMS08 0.162 0.228 $0.264$ $0.332$ $0.199$ $0.278$ $0.347$ $0.390$ $0.403$ $0.444$ $\underline{0.198}$ $\underline{0.236}$ $0.603$ $0.541$ $0.206$ $0.248$ $0.244$ $0.326$ $0.337$ $0.377$ $0.427$ $0.443$
Table 2: Short-term forecasting results. All the results are averaged from 4 different prediction lengths {12, 24, 48, 96}, and the look-back length is fixed to 96 for all baselines.
Adaptive Scale Routing Mixture-of-Experts
To address the limitations of static fusion mechanisms in multiscale forecasting, we propose the Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) block. This dynamic prediction head models temporal dependencies at different horizons via hierarchical experts (global-shared and local-specialized experts) and adaptively weights scale-specific predictions based on temporal patterns.
Specifically, ASR-MoE comprises two expert groups to handle distinct temporal granularities. Global experts $\mathcal{E}^{g}=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{m}\}$ capture common long-term dependencies through deep nonlinear transformations, modeling persistent trends and long-periodic patterns that commonly existed in time series. Local experts $\mathcal{E}^{l}=\{\mathcal{L}_{1},\mathcal{L}_{2},...,\mathcal{L}_{n}\}$ specialize in different short-term variations through shallower networks, detecting short-periodic patterns and high-frequency fluctuations that dynamically varied across stages. This explicit decoupling enables specialized handling of complex dynamics while maintaining higher parameter efficiency.
For weights calculation of each scale’s experts, a dynamic routing mechanism is proposed to assign input-dependent weights to experts, balancing global and local dependencies. Given a scale-specific feature $\mathbf{F}_{l}$, the router generates weights as follows:
$\mathbf{\Omega}_{l}=\mathrm{Softmax}(\mathrm{MLP}(\mathbf{F}_{l})),$ (10)
where the weight matrix $\mathbf{\Omega}$ are divided into global weights $\mathbf{\Omega}^{G}$ and local weights $\mathbf{\Omega}^{L}$. For computational efficiency, we activate only the top-$K$ local experts via sparse routing:
$\mathbf{\hat{\Omega}}^{L}_{l},\mathcal{I}=\mathrm{Top-K}(\mathbf{\Omega}^{L}_{l},K),$ (11)
where $\mathcal{I}$ denotes indices of selected experts.
Finally, a temporal-aware weighting module fuses outputs from all scales by historical scale importance. Scale-specific predictions $\hat{\mathbf{Y}}^{l}$ are fused using time-dependent weights:
$\mathbf{w}=\mathrm{Softmax}(\mathrm{MLP}(\bigoplus^{L}_{l=1}\phi(\mathbf{F}_{l})\cdot\mathbf{w}_{hist})),$ (12)
where $\phi(\cdot)$ denotes temporal descriptors, $\mathbf{w}_{hist}$ denotes historical weighting memory, and $\bigoplus$ means concatenation. The final prediction integrates multi-scale outputs:
$\hat{\mathbf{Y}}=\sum_{l=1}^{L}\mathbf{w}_{l}\left(\sum_{M}^{m=1}{\mathbf{\Omega}^{G}_{m}\mathcal{G}_{m}}(\mathbf{F}_{l})+\sum_{n\in{\mathcal{I}}}{\hat{\mathbf{\Omega}}^{L}_{n}\mathcal{L}_{n}}(\mathbf{F}_{l})\right),$ (13)
Meanwhile, to ensure balanced expert utilization, ASR-MoE incorporates a balance loss:
$\mathcal{L}_{{balance}}=-\lambda\mathbf{E}[\sum\mathbf{\Omega}_{j}log\mathbf{\Omega}_{j}],$ (14)
and the overall loss function of DMSC is as follows:
$\mathcal{L}=\mathcal{L}_{pred}+\mathcal{L}_{{balance}},$ (15)
where $\mathcal{L}_{pred}$ denotes the Mean Squared Error (MSE) loss.
By combining hierarchical expert specialization, dynamic routing, and temporal-aware weighting, ASR-MoE enables the proposed framework to adaptively prioritize relevant scales and experts, achieving a synergy between long-term trend projection and short-term detail refinement in TSF.
Experiments
To comprehensively verify the performance and effectiveness of DMSC, we conduct extensive experiments on 13 real-world TSF benchmarks spanning multiple domains and temporal resolutions.
Datasets.
For long-term forecasting, we conduct experiments on nine well-established benchmarks, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, ECL, Traffic, Weather and Solar. For short-term forecasting, we experiment on PEMS(PEMS03, PEMS04, PEMS07, PEMS08).
Baselines.
We compare our framework with ten SOTA models of TSF, including transformer-based models: iTransformer(Liu et al. 2024), PatchTST(Nie et al. 2023), Autoformer(Wu et al. 2021), TimeXer(Wang et al. 2024d); MLP-based models: DLinear(Zeng et al. 2023), PatchMLP(Tang and Zhang 2024), AMD(Hu et al. 2025); CNN-based models: TimesNet(Wu et al. 2023a), TimeMixer(Wang et al. 2024b), and a novel architecture TimeKAN(Huang et al. 2025).
Main Results
The main results are listed in Table 1 for long-term forecasting and Table 2 for short-term forecasting, in which the lower MSE and MAE indicates better forecasting performance. Our proposed DMSC framework demonstrates consistent SOTA performance across all 13 benchmarks, achieving the lowest MAE and MSE in most experimental configurations. These results confirm robust effectiveness and generality of DMSC for both long-term and short-term forecasting tasks. Notably, multi-scale architectures like TimesNet and PatchMLP exhibit instability due to their inflexible fusion mechanisms that fails to fully leverage captured multiscale features. And the preferable performance of TimeXer and iTransformer underscores the effect of the patch-wise and variate-wise strategy in information representation, but their fixed decomposition schemes inherently constrain the model’s potential. Compared to iTransformer and Dlinear, DMSC has a considerable improvement in data with more variate(Traffic, PEMS), demonstrating the necessity of jointly dependency modeling. Collectively, the above observations validate DMSC’s efficacy in addressing the core challenges of time series forecasting.
Model Analysis
Ablation Study.
[…]
Dateset Electricity Weather Traffic Solar
Metric MSE MAE MSE MAE MSE MAE MSE MAE
DMSC(Ours) 0.170 0.258 0.233 0.269 0.407 0.274 0.213 0.258
w/o EMPD $0.181$ $0.270$ $0.248$ $0.275$ $0.503$ $0.323$ $0.248$ $0.278$
w/o TIB $0.185$ $0.273$ $0.255$ $0.278$ $0.497$ $0.321$ $0.264$ $0.293$
w/o ASR-MoE $0.196$ $0.285$ $0.251$ $0.280$ $0.503$ $0.320$ $0.271$ $0.305$
Table 3: Ablation results. All results are averaged from prediction length {96, 192, 336, 720}.
Case Electricity Weather Traffic Solar
MSE MAE MSE MAE MSE MAE MSE MAE
DMSC(Ours) 0.170 0.258 0.233 0.269 0.407 0.274 0.213 0.258
[…]
Table 4: Ablation study on EMPD to compare dynamic decomposition with static predefined decomposition strategies.
Case Electricity Weather Traffic Solar
MSE MAE MSE MAE MSE MAE MSE MAE
DMSC(Ours) 0.170 0.258 0.233 0.269 0.407 0.274 0.213 0.258
[…]
Table 5: Ablation study on TIB to validate effectiveness of joint triad dependency modeling.
Case Electricity Weather Traffic Solar
MSE MAE MSE MAE MSE MAE MSE MAE
DMSC(Ours) 0.170 0.258 0.233 0.269 0.407 0.274 0.213 0.258
[…]
Table 6: Ablation study on ASR-MoE to assess the importance of expert specialization (global/local) and dynamic fusion strategies for prediction results.
For EMPD, employing predefined static patches or raw data embedding instead of input-adaptive decomposition degraded performance. This demonstrates the necessity of patch-wise decomposition and lightweight neural controllers for granularity adjustment. For TIB, replacing joint modeling with standard convolution layers or only intra-patch features resulted in performance drops, validating the need for comprehensive and joint modeling of triadic dependencies. Meanwhile, ablation of dynamic fusion mechanism in TIB shows its effectiveness to dynamically fuse these heterogeneous dependencies, as static handling of such relationships undermines holistic dependency capture across different scales. For ASR-MoE, substituting our adaptive prediction mechanism with either a single prediction head or summation of multiple linear heads yields substantial performance declines, which demonstrates that single heads inadequately utilize multiscale features, and simple aggregation of multiple heads overlooks the differential importance of scale-specific dependencies. Moreover, removing either global or local experts leads to performance degradation, which demonstrates the necessity of leveraging distinct experts for different temporal patterns. The elaborate design in ASR-MoE for collective utilization of specialized experts can effectively harness diverse temporal dynamics.
Collectively, these ablation studies confirm that dynamic and multiscale modeling is central to DMSC. All of three blocks and the proposed progressive cascade architecture together enable robust multiscale modeling and hierarchical feature learning. The consistent performance drops from ablating any component validate the integrity of DMSC’s architecture.
Figure 2: Visualization of different experts on ETTh1 datasets, the look-back and prediction length are set to 96.
ASR-MoE Expert Specialization.
To verify the specialization of ASR-MoE experts, we conducted activation pattern visualization across different experts. As shown in Fig. 2, distinct specialization profiles emerge: global experts exhibit pronounced sensitivity to persistent trends and long-periodic patterns, while local experts demonstrate acute responsiveness to short-periodic patterns and high-frequency fluctuations. This specialization demonstrates adaptively knowledge learning for different experts.
Increasing Look-Back Length.
Theoretically, longer input sequences provide richer historical information, potentially enhancing the model’s ability to capture long-term trends and complex multiscal dependencies. Our DMSC framework is specifically designed to leverage this, with its dynamic multiscale coordination capability inherently enabling adaptive modeling and effective utilization of information across varying look-back lengths. So we evaluate DMSC with look-back length from {48, 96, 192, 336, 720} on Electricity dataset. As shown in Fig. 3, DMSC consistently improves performance with longer historical sequences while outperforms other baselines, demonstrating its capacity to extract richer temporal representations from extended contexts and effectiveness in integrating multiscale temporal patterns.
Figure 3: Forecasting results with varying look-back length on Electricity dataset. Look-back lengths are set to {48, 96, 192, 336, 720}, and prediction length is fixed to 96.
Efficiency Analysis
We rigorously evaluate the memory usage and training time of DMSC against other SOTA baselines on Weather(21 variates) and Traffic datasets(862 variates). As demonstrated in Fig. 4 and Fig. 5, DMSC consistently achieves superior efficiency in both memory usage and training time while maintaining competitive forecasting accuracy. Crucially, DMSC demonstrates near-linear scalability to long sequences—significantly outperforming Transformer-based models that exhibit quadratic complexity and MLP-based approaches with parameter explosion. These efficiency gains are particularly pronounced when handling high-dimensional multivariate data (Traffic), where DMSC’s cross-variable interaction avoids the computational overhead of exhaustive pairwise attention while maintaining modeling capacity. Our framework thus establishes a balance of accuracy, latency, and memory usage for practical deployment scenarios.
Figure 4: Model efficiency analysis under 96-look-back length and 96-prediction length on Weather(21 variates) datasets. Batch size is set to 128.
Figure 5: Model efficiency analysis under 96-look-back length and 96-prediction length on Traffic(862 variates) datasets. Batch size is set to 16.
Conclusion
This paper introduces the Dynamic Multi-Scale Interaction (DMSC) framework, a novel framework that can advance TSF through dynamic multiscale coordination. DMSC has three key contributions: 1) an Embedded Multi-Scale Patch Decomposition(EMPD) is introduced to dynamically decompose time series into hierarchical patches to eliminate predefined and fixed scale constraints; 2) Triad Interaction Block (TIB) is proposed to jointly model intra-patch, inter-patch, and cross-variable dependencies, thus forming a coarse-to-fine feature pyramid through progressive cascade layers; 3) an Adaptive Scale Routing MoE (ASR-MoE) is proposed to dynamically fuse multiscale predictions via temporal-aware weighting of global and local experts. Extensive experiments on thirteen benchmarks demonstrate that DMSC can achieve SOTA performance and superior efficiency. Future work will extend the framework to multi-tasks learning and explore optimizations on intricate dataset environments, improving its multi-scenario applicability for real world.
Appendix A Implementation Details
Dataset Descriptions
We conduct long-term forecasting experiments on 9 real-world datasets, including ETT(ETTh1, ETTh2, ETTm1, ETTm2); Electricity(ECL); Exchange; Solar; Traffic; and Weather datasets. Meanwhile, we use PEMS(PEMS03, PEMS04, PEMS07, PEMS08) datasets to conduct short-term forecasting experiments. The details of datasets are listed in Tab 7.
Name Domain Length Num Prediction Length Dataset Size Freq. (m)
ETTh1 Temperature 14400 7 {96,192,336,720} (8545,2881,2881) 60
ETTh2 Temperature 14400 7 {96,192,336,720} (8545,2881,2881) 60
ETTm1 Temperature 57600 7 {96,192,336,720} (34465,11521,11521) 15
ETTm2 Temperature 57600 7 {96,192,336,720} (34465,11521,11521) 15
Electricity Electricity 26304 321 {96,192,336,720} (18317,2633,5261) 60
Exchange Exchange Rate 7588 8 {96,192,336,720} (5120,665,1422) 1440
Traffic Road Occupancy 17544 862 {96,192,336,720} (12185,1757,3509) 60
Weather Weather 52696 21 {96,192,336,720} (36792,5271,10540) 10
Solar-Energy Energy 52179 137 {96,192,336,720} (36601,5161,10417) 10
PEMS03 Traffic Flow 26208 358 {12,24,48,96} (15617,5135,5135) 5
PEMS04 Traffic Flow 16992 307 {12,24,48,96} (10172,3375,3375) 5
PEMS07 Traffic Flow 28224 883 {12,24,48,96} (16711,5622,5622) 5
PEMS08 Traffic Flow 17856 170 {12,24,48,96} (10690,3548,3548) 5
Table 7: Dataset Descriptions. Num is the number of variable. Dataset size is organized in (Train, Validation, Test).
Metrics Details
To evaluate model’s performance for TSF, we utilize the mean square error(MSE) and mean absolute error(MAE). The calculations of metrics are as follows:
$\textbf{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\mathbf{Y}_{i}-\hat{\mathbf{Y}}_{i})^{2}$ (16)
$\textbf{MAE}=\frac{1}{n}\sum_{i=1}^{n}|\mathbf{Y}_{i}-\hat{\mathbf{Y}}_{i}|$ (17)
Implementation Details
All experiments are implemented in Pytorch, and conducted on a single NVIDIA RTX 4090 24GB GPU. We utilize ADAM optimizer with an initial learning rate $10^{-3}$ and L2 loss for model optimization. For DMSC, we set layers of progressive cascade framework to 2 - 5, embedding dimension is set to 64, 128, 256, 512. The patch decay rate is set to 2 for exponentially degradation. All experiments are based on the framework of TimesNet. And all the baselines are implemented based on the configurations of original paper and its code.
Appendix B Model Analysis
Model Complexity Analysis
[…]
Hyperparameter Sensitivity Analysis
Embedding Dimension.
The embedding dimension determines the richness of feature representations. We set $d_{\mathrm{model}}\in\{64,128,256,512\}$ to evaluate model’s performance on ECL and Weather dataset. The results are shown in Fig. 6(a) and Fig. 6(b), we choose the optimal dimension at 128 / 256 - smaller embedding impairs multiscale feature separation, while larger configuration yield marginal gains with higher memory cost, indicating dimensionality saturating.
(a) Performance on Electricity dataset.
(b) Performance on Weather dataset.
Figure 6: Hyperparameter sensitivity analysis under different embedding dimension, $d_{\mathrm{model}}$ is set to {64, 128, 256, 512}. Left is conducted on Electricity dataset and right is on Weather dataset. Look-back length is fixed to 96, and prediction lengths are set to {96, 192, 336, 720}.
(a) Performance of different patch length decay rate.
(b) Performance of different progressive cascade layers.
Figure 7: Hyperparameter sensitivity analysis on Weather dataset. Left shows performance of different patch length decay rate, right shows performance of different progressive cascade layers. Look-back length is fixed to 96, and prediction lengths are set to {96, 192, 336, 720}.
EMPD Patch Length Decay.
EMPD employs exponentially decaying patch lengths to capture multiscale dependencies. We vary the decay rate (2, 3, 4) and find that a decay factor of 2 yields the best results from Fig. 7(a) on weather dataset. It can be seen that a decay factor of 2 yields the best forecasting performance, which balances the preservation of sufficient local details in fine-grained layers against the extraction of broader trends in coarse-grained layers.
Number of Progressive Cascade Layers.
The number of cascade layers ($l\in\{1-5\}$) affects the hierarchical feature extraction. As shown in Fig. 7(b), three layers can achieve peak efficiency - shallower stacks (1 - 2 layers) preserve perfect accuracy via residual propagation but miss fine-grained interactions, while deeper configurations (4 - 5 layers) introduce diminishing returns with much more latency but fewer improvement.
Appendix C Full Results
We evaluate DMSC on 13 real-world TSF benchmarks spanning diverse domains. Tab 8 shows the full results of long-term forecasting tasks on ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, ECL, Traffic, Weather and Solar datasets. Tab 9 contains the full results of short-term forecasting tasks on PEMS(PEMS03, PEMS04, PEMS07, PEMS08) datasets. From full results, we can see that DMSC achieves pretty good performance on most datasets and tasks.
Meanwhile, Tab 10 shows the full results of ablation study on Electricity, Weather, Traffic and Solar datasets, which demonstrating the efficiency and contribution of each component in DMSC.
Models DMSC(Ours) TimeMixer iTransformer PatchTST Dlinear TimesNet Autoformer TimeXer PatchMLP TimeKAN AMD
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.370 0.395 $0.379$ $\underline{0.397}$ $0.387$ $0.405$ $\underline{0.378}$ $0.399$ $0.397$ $0.412$ $0.415$ $0.429$ $0.589$ $0.526$ $0.386$ $0.404$ $0.393$ $0.405$ $0.387$ $0.401$ $0.394$ $0.404$
192 0.408 0.420 $0.430$ $0.429$ $0.441$ $0.436$ $0.427$ $0.429$ $0.446$ $0.441$ $0.479$ $0.466$ $0.653$ $0.551$ $0.429$ $0.435$ $0.443$ $0.434$ $\underline{0.415}$ $\underline{0.423}$ $0.444$ $0.432$
336 0.421 0.433 $0.493$ $0.459$ $0.494$ $0.463$ $0.468$ $0.455$ $0.489$ $0.467$ $0.517$ $0.482$ $0.715$ $0.581$ $0.484$ $0.457$ $0.486$ $0.456$ $\underline{0.453}$ $\underline{0.443}$ $0.485$ $0.451$
720 $\underline{0.462}$ 0.457 $0.522$ $0.493$ $0.488$ $0.483$ $0.508$ $0.497$ $0.513$ $0.511$ $0.505$ $0.490$ $0.726$ $0.601$ $0.544$ $0.513$ $0.509$ $0.487$ 0.461 $\underline{0.463}$ $0.486$ $0.472$
Avg 0.415 0.426 $0.456$ $0.444$ $0.452$ $0.446$ $0.445$ $0.445$ $0.461$ $0.457$ $0.479$ $0.466$ $0.670$ $0.564$ $0.460$ $0.452$ $0.458$ $0.445$ $\underline{0.429}$ $\underline{0.432}$ $0.452$ $0.439$
ETTh2
96 0.275 0.329 $0.290$ $0.341$ $0.301$ $0.350$ $0.295$ $0.347$ $0.341$ $0.394$ $0.316$ $0.358$ $0.443$ $0.459$ $\underline{0.284}$ $\underline{0.337}$ $0.311$ $0.358$ $0.291$ $0.342$ $0.397$ $0.451$
192 0.359 0.383 $\underline{0.366}$ $0.394$ $0.380$ $0.399$ $0.378$ $0.401$ $0.482$ $0.479$ $0.415$ $0.414$ $0.500$ $0.506$ $\underline{0.366}$ $\underline{0.391}$ $0.404$ $0.415$ $0.376$ $0.393$ $0.501$ $0.501$
336 0.376 0.398 $0.425$ $0.433$ $\underline{0.423}$ $\underline{0.431}$ $0.425$ $0.442$ $0.591$ $0.541$ $0.452$ $0.448$ $0.506$ $0.502$ $0.438$ $0.438$ $0.447$ $0.451$ $0.437$ $0.443$ $0.611$ $0.563$
720 $0.412$ 0.424 0.405 $\underline{0.431}$ $0.431$ $0.447$ $0.436$ $0.456$ $0.839$ $0.661$ $0.461$ $0.463$ $0.503$ $0.509$ $\underline{0.407}$ $0.449$ $0.462$ $0.464$ $0.451$ $0.458$ $0.956$ $0.718$
Avg 0.355 0.383 $\underline{0.372}$ $\underline{0.400}$ $0.384$ $0.407$ $0.383$ $0.412$ $0.563$ $0.519$ $0.411$ $0.421$ $0.488$ $0.494$ $0.374$ $0.404$ $0.406$ $0.422$ $0.389$ $0.409$ $0.616$ $0.558$
ETTm1
96 0.304 0.335 $0.319$ $0.359$ $0.341$ $0.376$ $0.326$ $0.366$ $0.346$ $0.374$ $0.336$ $0.375$ $0.564$ $0.506$ $\underline{0.318}$ $\underline{0.356}$ $0.321$ $0.362$ $0.326$ $0.365$ $0.331$ $0.363$
192 0.346 0.364 $\underline{0.360}$ $0.384$ $0.382$ $0.396$ $0.365$ $0.387$ $0.382$ $0.391$ $0.377$ $0.395$ $0.586$ $0.516$ $0.373$ $0.389$ $0.364$ $\underline{0.381}$ $\underline{0.360}$ $0.384$ $0.373$ $0.382$
336 $\underline{0.372}$ $\underline{0.392}$ $0.395$ $0.406$ $0.420$ $0.421$ $0.392$ $0.406$ $0.415$ $0.415$ $0.418$ $0.420$ $0.679$ $0.547$ $0.412$ 0.387 $0.396$ $0.404$ 0.369 $0.402$ $0.405$ $0.403$
720 $\underline{0.451}$ $\underline{0.440}$ $0.461$ $0.444$ $0.487$ $0.456$ $0.461$ $0.443$ $0.473$ $0.451$ $0.541$ $0.481$ $0.715$ $0.567$ $0.460$ $0.450$ $0.468$ $0.443$ 0.449 0.437 $0.467$ $0.437$
Avg 0.368 0.383 $0.384$ $0.398$ $0.408$ $0.412$ $0.386$ $0.400$ $0.404$ $0.408$ $0.418$ $0.418$ $0.636$ $0.534$ $0.391$ $0.395$ $0.387$ $0.398$ $\underline{0.376}$ $0.397$ $0.394$ $\underline{0.396}$
ETTm2
96 $\underline{0.174}$ $\underline{0.256}$ $0.179$ $0.261$ $0.186$ $0.272$ $0.184$ $0.269$ $0.193$ $0.293$ $0.188$ $0.268$ $0.554$ $0.469$ 0.172 0.254 $0.176$ $0.259$ $0.176$ $0.263$ $0.187$ $0.271$
192 0.233 0.295 $\underline{0.238}$ $\underline{0.301}$ $0.252$ $0.312$ $0.247$ $0.307$ $0.284$ $0.361$ $0.250$ $0.306$ $0.609$ $0.497$ $0.241$ $0.302$ $0.246$ $0.304$ $0.239$ $\underline{0.301}$ $0.251$ $0.309$
336 0.296 0.331 $\underline{0.299}$ $\underline{0.339}$ $0.315$ $0.351$ $0.313$ $0.354$ $0.382$ $0.429$ $0.306$ $0.341$ $0.401$ $0.409$ $0.301$ $0.340$ $0.309$ $0.344$ $0.304$ $0.346$ $0.309$ $0.346$
720 0.370 0.385 $0.395$ $\underline{0.394}$ $0.415$ $0.408$ $0.409$ $0.406$ $0.558$ $0.525$ $0.420$ $0.405$ $0.443$ $0.433$ $\underline{0.394}$ $0.395$ $0.417$ $0.408$ $0.410$ $0.408$ $0.407$ $0.400$
Avg 0.268 0.317 $0.278$ $0.324$ $0.292$ $0.336$ $0.288$ $0.334$ $0.354$ $0.402$ $0.291$ $0.330$ $0.502$ $0.452$ $\underline{0.277}$ $\underline{0.323}$ $0.287$ $0.329$ $0.282$ $0.330$ $0.288$ $0.332$
Electricity
96 0.138 0.223 $0.161$ $0.252$ $\underline{0.148}$ $\underline{0.241}$ $0.181$ $0.274$ $0.211$ $0.302$ $0.163$ $0.267$ $0.232$ $0.347$ $0.241$ $0.244$ $0.167$ $0.264$ $0.177$ $0.270$ $0.185$ $0.267$
192 $\underline{0.160}$ 0.258 $0.176$ $0.269$ $0.167$ $0.248$ $0.187$ $0.280$ $0.211$ $0.305$ $0.184$ $0.284$ $0.363$ $0.447$ 0.159 $\underline{0.260}$ $0.181$ $0.276$ $0.185$ $0.276$ $0.190$ $0.272$
336 0.167 0.253 $0.193$ $0.283$ $0.179$ $\underline{0.271}$ $0.204$ $0.296$ $0.223$ $0.319$ $0.196$ $0.297$ $0.599$ $0.595$ $\underline{0.177}$ $0.276$ $0.203$ $0.303$ $0.201$ $0.292$ $0.205$ $0.288$
720 $\underline{0.213}$ $\underline{0.299}$ $0.232$ $0.316$ 0.208 0.298 $0.246$ $0.328$ $0.258$ $0.351$ $0.232$ $0.325$ $0.775$ $0.701$ $0.229$ $0.321$ $0.251$ $0.341$ $0.241$ $0.323$ $0.246$ $0.321$
Avg 0.170 0.258 $0.190$ $0.280$ $\underline{0.176}$ $\underline{0.265}$ $0.205$ $0.295$ $0.226$ $0.319$ $0.194$ $0.293$ $0.492$ $0.523$ $0.201$ $0.275$ $0.200$ $0.296$ $0.201$ $0.290$ $0.206$ $0.287$
Exchange
96 0.082 0.201 $\underline{0.087}$ $\underline{0.204}$ $0.088$ $0.208$ $0.094$ $0.213$ $0.098$ $0.233$ $0.115$ $0.242$ $0.158$ $0.290$ $0.094$ $0.214$ $0.094$ $0.217$ $0.092$ $0.212$ $0.088$ $0.208$
192 0.174 $\underline{0.298}$ $\underline{0.177}$ 0.295 $0.180$ $0.303$ $0.182$ $0.303$ $0.186$ $0.325$ $0.216$ $0.333$ $0.299$ $0.406$ $0.182$ $0.303$ $0.187$ $0.311$ $0.180$ $0.300$ $0.182$ $0.305$
336 0.315 0.405 $0.328$ $\underline{0.414}$ $0.331$ $0.418$ $0.347$ $0.426$ $\underline{0.325}$ $0.434$ $0.375$ $0.444$ $0.470$ $0.511$ $0.384$ $0.448$ $0.342$ $0.424$ $0.352$ $0.430$ $0.332$ $0.417$
720 $\underline{0.772}$ 0.659 $0.847$ $0.692$ $0.848$ $0.695$ $0.931$ $0.724$ 0.746 $0.663$ $1.012$ $0.765$ $1.228$ $0.869$ $0.932$ $0.724$ $0.908$ $0.715$ $0.912$ $0.715$ $0.861$ $\underline{0.701}$
Avg 0.336 0.391 $0.359$ $\underline{0.401}$ $0.362$ $0.406$ $0.389$ $0.417$ $\underline{0.339}$ $0.414$ $0.430$ $0.446$ $0.539$ $0.519$ $0.398$ $0.422$ $0.383$ $0.417$ $0.384$ $0.414$ $0.366$ $0.408$
Weather
96 $\underline{0.160}$ $\underline{0.210}$ $0.164$ $\underline{0.210}$ $0.176$ $0.216$ $0.174$ $0.215$ $0.195$ $0.252$ $0.172$ $0.221$ $0.301$ $0.364$ 0.157 0.205 $0.168$ $0.214$ $0.163$ $0.210$ $0.194$ $0.236$
192 $\underline{0.207}$ $\underline{0.250}$ $0.208$ $0.251$ $0.225$ $0.257$ $0.221$ $0.256$ $0.239$ $0.299$ $0.220$ $0.260$ $0.335$ $0.385$ 0.204 0.248 $0.215$ $0.255$ $0.209$ $0.251$ $0.239$ $0.271$
336 0.253 0.284 $0.264$ $0.292$ $0.281$ $0.299$ $0.280$ $0.297$ $0.282$ $0.333$ $0.280$ $0.302$ $0.352$ $0.376$ $0.264$ $0.293$ $0.272$ $0.295$ $\underline{0.263}$ $\underline{0.291}$ $0.290$ $0.306$
720 0.313 0.332 $0.343$ $\underline{0.343}$ $0.361$ $0.353$ $0.357$ $0.349$ $0.345$ $0.381$ $0.353$ $0.350$ $0.405$ $0.401$ $0.343$ $\underline{0.343}$ $0.351$ $0.346$ $\underline{0.341}$ $0.344$ $0.362$ $0.352$
Avg 0.233 0.269 $0.245$ $0.274$ $0.261$ $0.281$ $0.258$ $0.279$ $0.265$ $0.316$ $0.256$ $0.283$ $0.348$ $0.382$ $\underline{0.242}$ $\underline{0.272}$ $0.252$ $0.277$ $0.244$ $0.274$ $0.271$ $0.291$
Traffic
96 0.389 0.259 $0.476$ $0.292$ $\underline{0.393}$ $\underline{0.269}$ $0.459$ $0.299$ $0.712$ $0.438$ $0.593$ $0.317$ $0.663$ $0.403$ $0.428$ $0.271$ $0.513$ $0.352$ $0.598$ $0.382$ $0.546$ $0.346$
192 0.405 0.258 $0.501$ $0.301$ $\underline{0.412}$ $\underline{0.277}$ $0.469$ $0.303$ $0.662$ $0.417$ $0.618$ $0.327$ $0.915$ $0.557$ $0.447$ $0.280$ $0.509$ $0.350$ $0.579$ $0.365$ $0.529$ $0.335$
336 0.398 $\underline{0.286}$ $0.514$ $0.314$ $\underline{0.424}$ 0.283 $0.483$ $0.309$ $0.669$ $0.419$ $0.642$ $0.341$ $1.217$ $0.704$ $0.472$ $0.289$ $0.533$ $0.360$ $0.572$ $0.361$ $0.540$ $0.339$
720 0.437 0.290 $0.545$ $0.320$ $\underline{0.459}$ $\underline{0.301}$ $0.518$ $0.326$ $0.709$ $0.437$ $0.679$ $0.350$ $1.317$ $0.755$ $0.517$ $0.307$ $0.599$ $0.395$ $0.609$ $0.384$ $0.576$ $0.358$
Avg 0.407 0.274 $0.509$ $0.307$ $\underline{0.422}$ $\underline{0.283}$ $0.482$ $0.309$ $0.688$ $0.428$ $0.633$ $0.334$ $1.028$ $0.605$ $0.466$ $0.287$ $0.539$ $0.364$ $0.590$ $0.373$ $0.547$ $0.345$
Solar
96 0.187 0.231 $\underline{0.196}$ $0.263$ $0.207$ $\underline{0.237}$ $0.216$ $0.274$ $0.290$ $0.378$ $0.223$ $0.256$ $0.552$ $0.524$ $0.198$ $0.244$ $0.239$ $0.272$ $0.248$ $0.302$ $0.310$ $0.312$
192 0.216 0.261 $0.245$ $0.279$ $0.242$ $\underline{0.264}$ $0.250$ $0.294$ $0.320$ $0.398$ $0.262$ $0.272$ $0.696$ $0.605$ $\underline{0.226}$ $0.270$ $0.291$ $0.298$ $0.291$ $0.321$ $0.348$ $0.330$
336 0.223 0.269 $\underline{0.235}$ $0.287$ $0.251$ $\underline{0.274}$ $0.265$ $0.302$ $0.353$ $0.415$ $0.287$ $0.299$ $0.816$ $0.677$ $0.239$ $0.281$ $0.325$ $0.316$ $0.307$ $0.332$ $0.399$ $0.355$
720 0.226 0.271 $\underline{0.236}$ $0.286$ $0.251$ $0.276$ $0.266$ $0.298$ $0.357$ $0.413$ $0.298$ $0.318$ $0.844$ $0.731$ $0.242$ $\underline{0.282}$ $0.336$ $0.321$ $0.310$ $0.329$ $0.393$ $0.349$
Avg 0.213 0.258 $0.228$ $0.279$ $0.238$ $\underline{0.263}$ $0.249$ $0.292$ $0.330$ $0.401$ $0.268$ $0.286$ $0.727$ $0.634$ $\underline{0.226}$ $0.269$ $0.298$ $0.301$ $0.289$ $0.321$ $0.363$ $0.337$
Table 8: Full results of long-term forecasting. All the results are selected from 4 different prediction lengths {96, 192, 336, 720}, and the look-back length is fixed to 96 for all baselines. A lower MSE or MAE indicates a better prediction, with the best in boldface and second in underline.
Models DMSC(Ours) TimeMixer iTransformer PatchTST Dlinear TimesNet Autoformer TimeXer PatchMLP TimeKAN AMD
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03
12 0.066 0.171 $0.084$ $0.194$ $0.069$ $\underline{0.175}$ $0.105$ $0.216$ $0.122$ $0.245$ $0.088$ $0.195$ $0.224$ $0.346$ $\underline{0.068}$ $0.179$ $0.118$ $0.191$ $0.095$ $0.210$ $0.106$ $0.225$
24 $\underline{0.092}$ 0.203 $0.130$ $0.244$ $0.099$ $0.210$ $0.198$ $0.296$ $0.202$ $0.320$ $0.118$ $0.224$ $0.492$ $0.513$ 0.089 $\underline{0.204}$ $0.204$ $0.231$ $0.166$ $0.281$ $0.174$ $0.294$
48 0.136 0.251 $0.218$ $0.317$ $0.164$ $0.275$ $0.472$ $0.466$ $0.334$ $0.428$ $0.169$ $0.268$ $0.392$ $0.459$ $\underline{0.137}$ $\underline{0.253}$ $0.213$ $0.314$ $0.314$ $0.393$ $0.333$ $0.417$
96 0.230 $\underline{0.339}$ $0.327$ $0.398$ $0.711$ $0.651$ $0.458$ $0.493$ $0.459$ $0.517$ $\underline{0.239}$ 0.330 $0.944$ $0.749$ $0.427$ $0.483$ $0.347$ $0.421$ $0.558$ $0.543$ $0.510$ $0.536$
Avg 0.131 0.241 $0.190$ $0.288$ $0.261$ $0.328$ $0.308$ $0.368$ $0.279$ $0.377$ $\underline{0.154}$ $\underline{0.254}$ $0.513$ $0.517$ $0.180$ $0.280$ $0.220$ $0.289$ $0.283$ $0.357$ $0.281$ $0.368$
PEMS04
12 0.078 0.183 $0.105$ $0.216$ $0.766$ $0.709$ $0.116$ $0.230$ $0.147$ $0.272$ $\underline{0.092}$ $\underline{0.202}$ $0.211$ $0.341$ $0.293$ $0.397$ $0.109$ $0.204$ $0.107$ $0.222$ $0.131$ $0.256$
24 0.102 0.215 $0.168$ $0.280$ $0.799$ $0.728$ $0.216$ $0.314$ $0.225$ $0.340$ $\underline{0.111}$ $\underline{0.224}$ $0.394$ $0.471$ $0.308$ $0.409$ $0.129$ $0.248$ $0.178$ $0.294$ $0.196$ $0.311$
48 0.147 0.261 $0.270$ $0.359$ $1.041$ $0.882$ $0.503$ $0.489$ $0.356$ $0.437$ $\underline{0.152}$ $\underline{0.266}$ $0.429$ $0.463$ $0.339$ $0.425$ $0.213$ $0.326$ $0.329$ $0.409$ $0.344$ $0.420$
96 0.190 $\underline{0.316}$ $0.377$ $0.439$ $1.045$ $0.886$ $0.623$ $0.586$ $0.453$ $0.505$ $\underline{0.197}$ 0.308 $0.853$ $0.703$ $0.367$ $0.441$ $0.361$ $0.441$ $0.572$ $0.563$ $0.638$ $0.602$
Avg 0.129 0.244 $0.230$ $0.324$ $0.913$ $0.801$ $0.365$ $0.405$ $0.295$ $0.389$ $\underline{0.138}$ $\underline{0.250}$ $0.472$ $0.495$ $0.327$ $0.418$ $0.203$ $0.305$ $0.297$ $0.372$ $0.327$ $0.397$
PEMS07
12 0.059 0.157 $0.070$ $0.173$ $0.068$ $0.169$ $0.093$ $0.206$ $0.116$ $0.241$ $0.075$ $0.179$ $0.207$ $0.335$ $\underline{0.061}$ $\underline{0.165}$ $0.107$ $0.178$ $0.085$ $0.199$ $0.096$ $0.223$
24 $\underline{0.078}$ $\underline{0.179}$ $0.109$ $0.215$ $0.087$ $0.190$ $0.195$ $0.295$ $0.209$ $0.327$ $0.083$ $0.198$ $0.314$ $0.412$ 0.071 0.177 $0.112$ $0.216$ $0.149$ $0.268$ $0.231$ $0.360$
48 $\underline{0.104}$ $\underline{0.211}$ $0.199$ $0.296$ $0.122$ $0.231$ $0.485$ $0.469$ $0.397$ $0.456$ $0.128$ $0.235$ $0.595$ $0.553$ 0.100 0.208 $0.168$ $0.282$ $0.292$ $0.383$ $0.143$ $0.272$
96 $\underline{0.128}$ $\underline{0.226}$ $0.283$ $0.341$ $0.159$ $0.267$ $0.979$ $0.716$ $0.592$ $0.552$ $0.150$ $0.253$ $0.556$ $0.563$ 0.120 0.221 $0.275$ $0.373$ $0.531$ $0.535$ $0.312$ $0.417$
Avg $\underline{0.092}$ $\underline{0.193}$ $0.165$ $0.256$ $0.109$ $0.214$ $0.438$ $0.422$ $0.329$ $0.394$ $0.109$ $0.216$ $0.418$ $0.466$ 0.088 0.192 $0.166$ $0.262$ $0.264$ $0.346$ $0.196$ $0.318$
PEMS08
12 0.076 0.178 $0.100$ $0.208$ $\underline{0.081}$ $\underline{0.183}$ $0.109$ $0.222$ $0.153$ $0.259$ $0.158$ $0.192$ $0.295$ $0.391$ $0.146$ $0.198$ $0.096$ $0.206$ $0.105$ $0.219$ $0.141$ $0.262$
24 $0.124$ $0.223$ $0.171$ $0.279$ $\underline{0.118}$ $\underline{0.222}$ $0.205$ $0.306$ $0.238$ $0.357$ 0.112 0.219 $0.345$ $0.419$ $0.171$ $0.221$ $0.144$ $0.256$ $0.179$ $0.291$ $0.243$ $0.350$
48 0.172 $\underline{0.216}$ $0.317$ $0.380$ $\underline{0.202}$ $0.292$ $0.493$ $0.484$ $0.473$ $0.515$ $0.231$ 0.198 $0.503$ $0.495$ $0.220$ $0.270$ $0.253$ $0.351$ $0.343$ $0.405$ $0.439$ $0.476$
96 0.278 0.296 $0.468$ $0.459$ $0.395$ $0.415$ $0.582$ $0.548$ $0.748$ $0.646$ $\underline{0.291}$ $0.338$ $1.268$ $0.857$ $0.285$ $\underline{0.303}$ $0.482$ $0.492$ $0.720$ $0.592$ $0.886$ $0.684$
Avg 0.162 0.228 $0.264$ $0.332$ $0.199$ $0.278$ $0.347$ $0.390$ $0.403$ $0.444$ $\underline{0.198}$ $\underline{0.236}$ $0.603$ $0.541$ $0.206$ $0.248$ $0.244$ $0.326$ $0.337$ $0.377$ $0.427$ $0.443$
Table 9: Full results of short-term forecasting. All the results are selected from 4 different prediction lengths {12, 24, 48, 96}, and the look-back length is fixed to 96 for all baselines. A lower MSE or MAE indicates a better prediction, with the best in boldface and second in underline.
[…]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity
96 0.138 0.223 $0.155$ $0.247$ $0.155$ $0.247$ $0.175$ $0.267$ $0.151$ $0.245$ $0.161$ $0.251$ $0.157$ $0.250$ $0.166$ $0.256$ $0.160$ $0.251$ $0.171$ $0.259$
192 0.160 0.258 $0.170$ $0.259$ $0.169$ $0.259$ $0.182$ $0.275$ $0.163$ $0.254$ $0.171$ $0.261$ $0.166$ $0.257$ $0.177$ $0.266$ $0.167$ $0.259$ $0.176$ $0.264$
336 0.167 0.253 $0.179$ $0.270$ $0.186$ $0.276$ $0.192$ $0.282$ $0.176$ $0.271$ $0.188$ $0.278$ $0.186$ $0.278$ $0.194$ $0.283$ $0.192$ $0.284$ $0.202$ $0.290$
720 0.213 0.299 $0.220$ $0.305$ $0.228$ $0.311$ $0.234$ $0.317$ $0.225$ $0.312$ $0.228$ $0.312$ $0.238$ $0.321$ $0.238$ $0.319$ $0.245$ $0.325$ $0.246$ $0.328$
Avg 0.170 0.258 $0.181$ $0.270$ $0.185$ $0.273$ $0.196$ $0.285$ $0.179$ $0.271$ $0.187$ $0.276$ $0.187$ $0.277$ $0.194$ $0.281$ $0.191$ $0.280$ $0.199$ $0.285$
Weather
96 0.160 0.210 $0.161$ $0.208$ $0.173$ $0.216$ $0.167$ $0.215$ $0.160$ $0.205$ $0.166$ $0.210$ $0.167$ $0.216$ $0.173$ $0.219$ $0.170$ $0.219$ $0.164$ $0.213$
192 0.207 0.250 $0.213$ $0.253$ $0.218$ $0.255$ $0.216$ $0.260$ $0.215$ $0.253$ $0.214$ $0.254$ $0.213$ $0.254$ $0.222$ $0.260$ $0.213$ $0.254$ $0.213$ $0.252$
336 0.253 0.284 $0.269$ $0.292$ $0.274$ $0.296$ $0.271$ $0.297$ $0.264$ $0.289$ $0.269$ $0.294$ $0.266$ $0.291$ $0.271$ $0.296$ $0.273$ $0.296$ $0.256$ $0.287$
720 0.313 0.332 $0.349$ $0.346$ $0.353$ $0.343$ $0.348$ $0.349$ $0.342$ $0.340$ $0.346$ $0.343$ $0.343$ $0.332$ $0.352$ $0.348$ $0.346$ $0.344$ $0.338$ $0.331$
Avg 0.233 0.269 $0.248$ $0.275$ $0.255$ $0.278$ $0.251$ $0.280$ $0.245$ $0.272$ $0.249$ $0.275$ $0.247$ $0.273$ $0.255$ $0.281$ $0.251$ $0.278$ $0.243$ $0.271$
Traffic
96 0.389 0.259 $0.479$ $0.311$ $0.486$ $0.318$ $0.484$ $0.314$ $0.443$ $0.299$ $0.461$ $0.303$ $0.459$ $0.308$ $0.494$ $0.301$ $0.494$ $0.320$ $0.451$ $0.317$
192 0.405 0.258 $0.493$ $0.316$ $0.483$ $0.314$ $0.491$ $0.313$ $0.458$ $0.300$ $0.484$ $0.311$ $0.473$ $0.309$ $0.495$ $0.308$ $0.487$ $0.317$ $0.486$ $0.329$
336 0.398 0.286 $0.503$ $0.322$ $0.490$ $0.315$ $0.503$ $0.318$ $0.471$ $0.302$ $0.493$ $0.318$ $0.485$ $0.311$ $0.507$ $0.316$ $0.505$ $0.326$ $0.479$ $0.317$
720 0.437 0.290 $0.538$ $0.341$ $0.530$ $0.337$ $0.535$ $0.334$ $0.542$ $0.349$ $0.516$ $0.327$ $0.516$ $0.328$ $0.544$ $0.321$ $0.531$ $0.334$ $0.535$ $0.342$
Avg 0.407 0.274 $0.503$ $0.323$ $0.497$ $0.321$ $0.503$ $0.320$ $0.479$ $0.313$ $0.489$ $0.315$ $0.483$ $0.314$ $0.510$ $0.312$ $0.504$ $0.324$ $0.488$ $0.326$
Solar
96 0.187 0.231 $0.221$ $0.260$ $0.232$ $0.278$ $0.242$ $0.290$ $0.208$ $0.255$ $0.237$ $0.280$ $0.240$ $0.278$ $0.241$ $0.284$ $0.226$ $0.272$ $0.234$ $0.279$
192 0.216 0.261 $0.241$ $0.273$ $0.264$ $0.295$ $0.295$ $0.324$ $0.248$ $0.281$ $0.281$ $0.298$ $0.302$ $0.316$ $0.283$ $0.303$ $0.279$ $0.304$ $0.273$ $0.303$
336 0.223 0.269 $0.264$ $0.290$ $0.283$ $0.306$ $0.266$ $0.302$ $0.261$ $0.289$ $0.289$ $0.309$ $0.301$ $0.312$ $0.289$ $0.301$ $0.276$ $0.296$ $0.305$ $0.322$
720 0.226 0.271 $0.267$ $0.288$ $0.276$ $0.293$ $0.282$ $0.304$ $0.254$ $0.280$ $0.292$ $0.303$ $0.264$ $0.289$ $0.298$ $0.306$ $0.280$ $0.299$ $0.289$ $0.306$
Avg 0.213 0.258 $0.248$ $0.278$ $0.264$ $0.293$ $0.271$ $0.305$ $0.243$ $0.276$ $0.275$ $0.298$ $0.277$ $0.299$ $0.278$ $0.299$ $0.265$ $0.293$ $0.275$ $0.303$
Table 10: Full results of ablation study. All the results are selected from 4 different prediction lengths {96, 192, 336, 720}, and the look-back length is fixed to 96 for all baselines. A lower MSE or MAE indicates a better prediction, with the best in boldface.

### v4

@@
 Comprehensive experiments on thirteen real-world benchmarks demonstrate that DMSC consistently maintains state-of-the-art (SOTA) performance and superior computational efficiency for TSF tasks.
-Code is available at https://github.com/1327679995/DMSC.
 Introduction
@@
 2020; Jin et al.
-2021), weather forecasting(Wu et al.
-2023b; Bi et al.
-2023), economics(Chen et al.
+2021), weather forecasting (Bi et al.
+2023; Wu et al.
+2023b) and economics (Chen et al.
 2023b; Yu et al.
-2023) and urban computing(Li et al.
-2024; Yuan et al.
+2023).
+Time series data exhibit intricate temporal dependencies, which include nonlinear relationships, multi-scale patterns (e.g., trends and seasonality), and dynamic variable couplings (Qiu et al.
 2024).
-Time series data exhibit intricate temporal dependencies, encompassing nonlinear relationships, long-term and short-term patterns such as trends, seasonality and dynamic variable couplings(Qiu et al.
-2024).
-Owing to non-stationarity(Nason 2006) and complex interconnections(Dagum and Bianconcini 2016), traditional approaches struggle to effectively deal with time series data, making effectively dependency modeling a critical challenge for enhancing predictive precision(Shao et al.
+Due to the inherent non-stationarity (Nason 2006) and complex interconnections (Dagum and Bianconcini 2016), traditional approaches often fail to adequately capture the underlying patterns in time series data, making effectively dependency modeling a critical challenge for enhancing predictive accuracy (Shao et al.
 2025).
-In recent years, the powerful ability of deep learning to capture complex dependency relationships in time series data has been demonstrated, making it a formidable tool in the field of TSF(Chen et al.
+In recent years, deep learning has demonstrated a strong capability to capture complex dependency relationships in time series data, making it a formidable tool for TSF (Chen et al.
 2023a).
-However, intricate dependencies in time series inherently manifest across multiscales: coarser granularities encapsulate long-term trends while finer resolutions capture short-term fluctuations and periodic patterns.
-Existing single-scale modeling approaches often fail to balance fine-grained local details and holistic global trends(Wang et al.
+However, intricate dependencies in time series inherently manifest across multiscales: coarser granularities typically encapsulate long-term trends while finer resolutions capture short-term fluctuations and periodic patterns.
+Conventional single-scale modeling approaches often fail to balance fine-grained local details and holistic global trends (Wang et al.
 2025).
-To effectively model multiscale temporal dependencies, recent research has pioneered multiscale modeling frameworks that concurrently extract global and local patterns, thereby mitigating the representational limitations of single-scale approaches.
+To effectively model multi-scale temporal dependencies, recent research has pioneered multi-scale modeling frameworks that concurrently extract global and local patterns, thereby thereby mitigating the representational constraints inherent in single-scale approaches.
 While previous implementations often employ moving average downsampling (Wang et al.
 2024b; Huang et al.
-2025), which often remains insufficient in capturing local semantic information.
+2025), which remains insufficient in capturing local semantic information.
 In contrast, more effective alternatives focus on patch-wise and variable-wise decomposition operations that decompose sequence across feature and temporal dimensions.
-Such decomposition operations better preserves both global trends and local semantics, and multi-length patches inherently learn multi-scale representations.
-Critically, prevailing methods exhibit two constraints:1) dependence on fixed-scale decomposition strategies that lack dynamic scale adaptation; 2) fragmented processing of temporal dependencies and cross-variable interactions, undermining holistic dependency modeling.
-The core value of multiscale feature extraction resides in mining complementary information across diverse temporal granularities and hierarchical levels, thereby overcoming the limitations of single-scale modeling.
-Following feature capture, the effective design of prediction heads for multiscale fusion is equally important.
-Conventional approaches typically employ either linear projection layers(Li, Li, and Diao 2025), additive combinations of multiple projections(Wang et al.
+Such decomposition operations better preserve both global trends and local semantics, and multi-length patches inherently facilitate the learning of multi-scale representations.
+Critically, prevailing methods suffer from two limitations: 1) reliance on fixed-scale decomposition strategies that lack dynamic scale adaptation; 2) fragmented processing of temporal dependencies and cross-variable interactions, which compromises holistic dependency modeling.
+The primary advantage of multi-scale feature extraction lies in its capacity to mine complementary information across diverse temporal granularities and hierarchical levels, thereby overcoming the limitations of single-scale modeling.
+Following feature capture, the effective design of prediction heads for multi-scale fusion is equally important.
+Conventional approaches typically employ linear projection layers (Li, Li, and Diao 2025), additive combinations of multiple projections (Wang et al.
 2024b), or spectral amplitude-weighted aggregation (Wu et al.
 2023a).
-However, above simplistic fusion method exhibit fundamental inadequacies: they fail to leverage heterogeneous dominance patterns across multiscales, disregard dynamic inter-scale importance, and induce quadratic parameter growth with increasing input lengths and scale counts, severely compromising deployment efficiency(Qiu et al.
+However, these simplistic fusion methods exhibit fundamental inadequacies: they fail to leverage heterogeneous dominance patterns across multiscales, disregard dynamic inter-scale importance, and incur quadratic growth of model parameters with increasing input lengths and scale counts, thereby severely compromising deployment efficiency (Qiu et al.
 2024).
@@
 It effectively captures intricate dependencies across temporal resolutions and cross-variable interactions.
-Meanwhile, it can adaptively predict learned representations through temporal-aware fusion.
-Specifically, DMSC processes intricate dependencies through a multi-layer progressive architecture and Adaptive Scale Routing Mixture-of-Experts (ASR-MoE).
-The multi-layer progressive cascade architecture incorporates Embedded Multi-Scale Patch Decomposition Block (EMPD), Triad Interaction Block (TIB), where EMPD transforms input into 3D representations at varying granularities and TIB models heterogeneous dependencies within each layer.
+Meanwhile, it can generate adaptive predictions based on learned representations via temporal-aware fusion.
+Specifically, DMSC processes intricate dependencies through a multi-layer progressive architecture and an Adaptive Scale Routing Mixture-of-Experts (ASR-MoE).
+The multi-layer progressive cascade architecture incorporates Embedded Multi-Scale Patch Decomposition Block (EMPD), Triad Interaction Block (TIB), in which EMPD transforms the input into 3D representations at varying granularities and TIB models heterogeneous dependencies within each layer.
 In this architecture, coarse-grained representations from shallow layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways.
-After multi-scale feature extraction, ASR-MoE is employed to perform time-aware weighted dynamic fusion for multi-scale prediction, enabling adaptive integration of features across scales.
-By synergistically integrating dynamic patch decomposition, deeper models and sparse MoE principles, DMSC achieves full-spectrum multi-scale coordination across embedding, feature extraction, and prediction stages.
-This dynamic co-optimization enable it to achieve SOTA performance across 13 real-world benchmarks while maintaining high efficiency and low computational complexity.
+After multi-scale feature extraction, ASR-MoE performs temporal-aware weighted dynamic fusion for multi-scale prediction, thus enabling adaptive integration of cross-scale features to enhancing forecasting accuracy.
+DMSC formulates a unified framework that incorporates dynamic patch decomposition, deeper models, and sparse MoE principles to achieve full-spectrum multi-scale coordination.
+This dynamic synergy enables the DMSC framework to achieve SOTA performance across 13 real-world benchmarks while maintaining high efficiency and low computational cost.
 Our contributions are as follows:
   • A novel multi-layer progressive cascade architecture is designed to jointly integrate Embedded Multi-Scale Patch Decomposition Block (EMPD) and Triad Interaction Block (TIB) across layers.
-Unlike fixed-scale approaches, EMPD dynamically adjusts patch granularities by employing a lightweight neural controller based on input characteristics, and TIB jointly models intra-patch, inter-patch, and inter-variable dependencies through gated feature fusion to form a coarse-to-fine feature pyramid, where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated residual pathways.
+Unlike fixed-scale approaches, EMPD dynamically adjusts patch granularities by employing a lightweight network based on the temporal characteristics of input data, and TIB jointly models intra-patch, inter-patch, and inter-variable dependencies through gated feature fusion to form a coarse-to-fine feature pyramid, where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated residual pathways.
   • Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) is proposed to resolve static fusion limitations in prediction.
-ASR-MoE establishes a hierarchical expert architecture to explicitly decouples long-term and short-term dependencies, where global-shared experts capture common long-term temporal dependencies, while local-specialized experts model different short-term variations.
+It establishes a hierarchical expert architecture to explicitly decouples long-term and short-term dependencies, in which global-shared experts capture common long-term temporal dependencies, while local-specialized experts model different short-term variations.
 Significantly, a temporal-aware weighting aggregator is designed to dynamically compute scale-specific prediction contributions with historical memory.
@@
 2024c).
-MLP-based models(Oreshkin et al.
+MLP-based models (Chi et al.
+2021; Oreshkin et al.
 2020; Yi et al.
-2023; Chi et al.
-2021) primarily model temporal dependencies through predefined decomposition schemes and fully-connected architectures.
-While demonstrating simplicity and efficiency in capturing intra-series patterns, they exhibit critical limitations in modeling long-term dependencies and complex inter-series relationships.
+2023) typically capture temporal dependencies through predefined decomposition and fully-connected layers.
+While these models demonstrate efficiency in capturing intra-series patterns, they exhibit notable limitations in modeling long-term dependencies and complex inter-series relationships.
 In contrast, CNN-based models (Bai, Kolter, and Koltun 2018; LIU et al.
@@
 2023a).
-Nevertheless, Transformer-based models(Zhou et al.
-2021; Wu et al.
-2021; Wang et al.
-2024a) have emerged as a powerful paradigm by utilizing self-attention mechanisms(Vaswani et al.
-2017) to model superior long-term dependency across long sequences, so that effectively capturing persistent temporal relationships.
-However, Transformer-based models face substantive criticism regarding permutation invariance and quadratic computational complexity limitations in recent researches(Zeng et al.
-2023; Tang and Zhang 2024).
-While significant progress has been made, existing deep architectures primarily operate on single or fixed decomposition, and often model dependencies in a fragmented manner, lacking a mechanism for dynamic, adaptive multi-scale decomposition and coordinated dependency modeling.
+Nevertheless, Transformer-based models (Wang et al.
+2024a; Wu et al.
+2021; Zhou et al.
+2021) have emerged as a powerful paradigm, leveraging self-attention mechanisms (Vaswani et al.
+2017) to effectively capture long-range dependencies and persistent temporal relationships.
+However, Transformer-based models face substantial criticism due to their permutation invariance and quadratic computational complexity limitations in recent researches (Tang and Zhang 2024; Zeng et al.
+2023).
+Despite significant progress, most existing deep architectures rely on single or fixed decomposition strategies, and often model dependencies in a fragmented manner, lacking mechanisms for dynamic, adaptive multi-scale decomposition and coordinated dependency modeling.
 Decomposition Operations in TSF
-To effectively address intricate temporal dependencies, most studies have adopted multiscale modeling perspective, employing diverse multiscale decomposition operations on time series.
+To effectively capture intricate temporal dependencies, numerous studies have adopted a multi-scale modeling perspective, employing diverse decomposition operations on time series.
 TimesNet (Wu et al.
-2023a) learns frequency-adaptive periods in the frequency domain, transforming time series into 2D tensors to explicitly model intra-period and inter-period variations.
+2023a) learns frequency-adaptive periods in the frequency domain and transforms series into 2D tensors, explicitly modeling intra-period and inter-period variations.
 TimeMixer (Wang et al.
 2024b) decomposes sequences into seasonal and trend components (Wu et al.
-2021), subsequently applying different granularity mixing strategies to integrate multiscale features.
+2021), then applies mixing strategies at different granularities to integrate multi-scale features.
 However, simply using moving average inadequately preserves local semantic patterns.
-Diverging from the above operations, some methods focus on variable-wise and patch-wise disentanglement.
+Diverging from these operations, several methods focus on variable-wise and patch-wise disentanglement.
 PatchTST (Nie et al.
-2023) segments series into subseries-level patches as input tokens, employing channel independence to capture local semantics.
+2023) segments series into subseries-level patches as input tokens and employs channel independence to capture local semantics.
 ITransformer (Liu et al.
-2024) embeds entire series as singular tokens to model extended global representations.
+2024) embeds entire series as single tokens to model extended global representations.
 TimeXer (Wang et al.
 2024d) hierarchically represents variables through dual disentanglement, variable-wise for cross-channel interactions and patch-wise for endogenous variables.
-However, these decomposition techniques are constrained by static or predefined scale configurations, which hinders their adaptability to complex temporal patterns.
-Figure 1: The framework of DMSC. EMPD dynamically segments input series into hierarchical patches, TIB jointly models three dependencies, and ASR-MoE fuses multiscale predictions adaptively.
-Multiscale Prediction Operations in TSF
-Meanwhile, researches also focuses on multiscale operations in prediction heads.
+However, these decomposition techniques are constrained by static or predefined scale settings, hindering their adaptability to complex temporal patterns.
+Figure 1: The framework of DMSC. EMPD dynamically segments input series into hierarchical patches, TIB jointly models three dependencies, and ASR-MoE fuses multi-scale predictions adaptively.
+Multi-Scale Prediction Operations in TSF
+Beyond multi-scale decomposition operations, researchers have also investigated multi-scale operations in prediction heads.
 TimeMixer (Wang et al.
-2024b) employs dedicated projectors-one per scale-to generate scale-specific predictions.
+2024b) employs dedicated projectors to generate scale-specific predictions.
 Diverging from simplistic additive aggregation, TimeMixer++ (Wang et al.
-2025) utilizes frequency-domain amplitude analysis to weightedly aggregate projector outputs.
+2025) utilizes frequency-domain amplitude analysis to perform weighted aggregation of projector outputs.
 TimeKAN (Huang et al.
-2025) adopts frequency transformation and padding operations to unify multiscale representations into identical dimensions, enabling holistic scale integration.
-Current multiscale fusion mechanisms predominantly rely on static weighting schemes or fixed designs.
-They fail to dynamically prioritize different scales with intricate temproal dependencies.
-In contrast to these fragmented approaches, DMSC pioneers fully-spectrum multiscale coordination spanning embedding, extraction, and prediction stages, achieving optimal balance between computational efficiency and forecasting accuracy.
+2025) adopts frequency transformation and padding operations to unify multi-scale representations into identical dimensions, thereby enabling holistic scale integration.
+Current multi-scale fusion mechanisms predominantly rely on static weighting or fixed designs.
+They fail to dynamically prioritize the importance of different scales with intricate temporal dependencies.
+Different from these fragmented approaches, our proposed DMSC framework explores full-spectrum multi-scale coordination across embedding, extraction, and prediction stages, achieving both superior forecasting accuracy and high computational efficiency.
 Dynamic Multi-Scale Coordination Framework
-Time series forecasting addresses the fundamental challenge of predicting future values ${Y}={y_{t+1},...,y_{t+\tau}}\in\mathbf{R}^{\tau\times{C}}$ from historical observations ${X}={x_{1},...,x_{t}}\in\mathbf{R}^{t\times{C}}$, where $\tau$ is the prediction horizon, $t$ is the history horizon, and $C$ is the number of variate.
-To model complex multiscale dependencies inherent in real-world series, the Dynamic Multi-Scale Coordination (DMSC) framework is proposed.
+Time series forecasting addresses the fundamental challenge of predicting future values ${Y}={y_{t+1},...,y_{t+h}}\in\mathbb{R}^{h\times{C}}$ from historical observations ${X}={x_{1},...,x_{t}}\in\mathbb{R}^{t\times{C}}$, where $h$ is the prediction horizon, $t$ is the history horizon, and $C$ is the number of variables.
+To model complex multi-scale dependencies inherent in real-world time series, we propose the Dynamic Multi-Scale Coordination (DMSC) framework.
 As illustrated in Fig.
-1, DMSC establishes full-spectrum multi-scale coordination across embedding, feature extraction, and prediction stages using EMPD, TIB and ASR-MoE.
+1, DMSC establishes full-spectrum multi-scale coordination across three key stages: embedding, feature extraction, and prediction, realized through EMPD, TIB and ASR-MoE components.
 Multi-Layer Progressive Cascade Architecture
-The multi-layer progressive cascade architecture enables hierarchical feature learning through stacked EMPD-TIB units, creating a dynamic feature pyramid where coarse-grained representations from earlier layers adaptively guide fine-grained extraction in subsequent layers via gated pathways.
-Formally, given input $\mathbf{X}\in\mathbf{R}^{C\times L}$, layer $l$ processes:
+The multi-layer progressive cascade architecture facilitates hierarchical feature learning through stacked EMPD-TIB units, forming a dynamic feature pyramid.
+Within this pyramid, coarse-grained representations from earlier layers adaptively guide fine-grained feature extraction in subsequent layers via gated pathways.
+Formally, given input $\mathbf{X}\in\mathbb{R}^{C\times L}$, the $l$-th layer processes:
 $\mathcal{T}=\{\mathbf{F}_{1},\mathbf{F}_{2},...,\mathbf{F}_{l}\},$
@@
 $\mathbf{Z}_{l}=\textbf{EMPD}_{l}(\mathcal{G}_{l}(\mathbf{F}_{l-1})+\mathbf{X}),$ (1)
-where $\mathcal{T}$ is the set of hierarchical features generated by this architecture for ASR-MoE, with each $\mathbf{F}_{l}\in\mathbf{R}^{C\times{D}}$ representing the $l$-th layer’s output.
-$\mathcal{G}_{l}$ is a gated projection matrix that adaptively modulates the residual flow.
-This cascade flow enables progressive refinement of multiscale representations, where $\mathbf{F}_{l-1}$ adaptively guides $\mathbf{Z}_{l}$ with residual connections, and $\textbf{TIB}_{l}$ iteratively enhances representations by jointly integrating intra-patch, inter-patch, and cross-variable dependencies.
-By interleaving dynamic patch decomposition (EMPD) with triadic dependency modeling (TIB), our cascade architecture achieves input-aware progressive refinement of temporal granularities, enabling coherent modeling of both short-term dynamics and long-range trends while maintaining parameter efficiency.
+where $\mathcal{T}$ denotes the set of hierarchical features generated by this architecture for ASR-MoE, with each $\mathbf{F}_{l}\in\mathbb{R}^{C\times{D}}$ representing the output of $l$-th layer.
+Here, $\mathcal{G}_{l}$ represents a gated projection matrix that adaptively modulates the residual information flow.
+This cascade flow enables progressive refinement of multi-scale representations, where $\mathbf{F}_{l-1}$ adaptively guides the computation of $\mathbf{Z}_{l}$ with residual connections, and $\textbf{TIB}_{l}$ iteratively enhances representations by jointly modeling intra-patch, inter-patch, and cross-variable dependencies.
+By combining dynamic patch decomposition (EMPD) with triadic dependency modeling (TIB), our cascade architecture achieves input-aware progressive refinement of temporal granularities.
+This integration enables coherent modeling of both short-term dynamics and long-range trends while maintaining parameter efficiency.
 Embedded Multi-Scale Patch Decomposition
-Patch decomposition serves as a fundamental operation to transform time series into structured representations that preserve local semantics while enabling hierarchical pattern discovery.
-Unlike conventional single-fixed or fixed-scale approaches, the EMPD block introduces input-adaptive and hierarchical decomposition mechanism as a built-in component, dynamically adjusting patch granularities based on sequence characteristics.
-Thus eliminating rigid predefined patch configurations while natively integrating multiscale decomposition into the model architecture.
-Specifically, given an input time series tensor $\mathbf{X}\in\mathbf{R}^{C\times{L}}$(where $C$ is the number of features, $L$ is the sequence length), EMPD first computes a scale factor $\alpha$ through a lightweight neural controller:
+Patch decomposition serves as a fundamental operation that transforms time series into structured representations, preserving local semantics while enabling hierarchical pattern discovery.
+Unlike conventional single or fixed-scale approaches, the EMPD block introduces adaptive and hierarchical decomposition mechanism as a built-in component, dynamically adjusting patch granularities based on sequence characteristics.
+This approach eliminates rigid predefined patch configurations while natively integrating multi-scale decomposition into our framework.
+Specifically, given an input time series tensor $\mathbf{X}\in\mathbb{R}^{C\times{L}}$(where $C$ denotes the number of variates and $L$ the sequence length), EMPD first computes a scale factor $\alpha$ through a lightweight neural network:
 $\alpha=\mathcal{N}_{\theta}(\Phi_{\mathrm{GAP}}(\mathbf{X}))$ (2)
-where $\mathcal{N}_{\theta}$ is a micro-MLP with sigmoid activation, $\Phi_{\mathrm{GAP}}$ compresses temporal dimensions through averaging.
-Factor $\alpha\in[0,1]$ dynamically determines the base patch length, and then the patch length decays exponentially as follows:
-$\mathbf{P}_{{base}}=[\mathbf{P}_{min}+\alpha\cdot(\mathbf{P}_{max}-\mathbf{P}_{min})],$
-$\mathbf{P}_{l}=\mathrm{max}(\mathbf{P}_{min},[\mathbf{P}_{base}/\tau^{l}]),$ (3)
-where $\mathbf{P}_{max}$ and $\mathbf{P}_{min}$ denoting the learned parameter for minimum and maximum patch bounds, and $\mathbf{P}_{l}$ means patch length in layer $l$, which ensures shallow layers process coarse-grained dependencies while deeper layers capture fine-grained patterns.
-Then EMPD applies replication padding to handle boundary effects, followed by hierarchical patch unfolding.
-The operations are formalized as:
-$\mathbf{X}_{p}^{l}=\mathrm{Unfold}(\mathrm{Padding}_{S_{l}}(X)),$ (4)
-where $\mathrm{Padding}_{S_{l}}(\cdot)$ denotes replication padding with stride $S_{l}=P_{l}/2$ and $\mathrm{Unfold}_{P_{l},S_{l}}$ unfolds the padded sequence into patches of length $P_{l}$ with stride $S_{l}$.
-This operation yields a 3D patch tensor $\mathbf{X}_{u}^{l}\in\mathbf{R}^{C\times{N_{l}}\times{P_{l}}}$.
-Then EMPD transforms $\mathbf{X}_{u}^{l}$ into a unified embedding space via linear projection:
+where $\mathcal{N}_{\theta}$ denotes a lightweight MLP with sigmoid activation, and $\Phi_{\mathrm{GAP}}$ denotes a global average pooling operation that compresses the temporal dimension.
+The factor $\alpha\in[0,1]$ dynamically determines the base patch length, which subsequently undergoes exponential decay across layers as follows:
+$\mathbf{P}_{\text{base}}=[\mathbf{P}_{\text{min}}+\alpha\cdot(\mathbf{P}_{\text{max}}-\mathbf{P}_{\text{min}})],$
+$\mathbf{P}_{l}=\mathrm{max}(\mathbf{P}_{\text{min}},[\mathbf{P}_{\text{base}}/\tau^{l}]),$ (3)
+where $\mathbf{P}_{\text{max}}$ and $\mathbf{P}_{\text{min}}$ denote the parameters for minimum and maximum patch bounds, respectively.
+$\mathbf{P}_{l}$ represents the patch length at the $l$-th layer.
+This design ensures that shallow layers process coarse-grained dependencies while deeper layers capture fine-grained ones.
+EMPD then applies replication padding to mitigate boundary effects, followed by hierarchical patch unfolding operation:
+$\mathbf{X}_{p}^{l}=\mathrm{Unfold}_{P_{l},S_{l}}(\mathrm{Padding}_{S_{l}}(X)),$ (4)
+where $\mathrm{Padding}_{S_{l}}(\cdot)$ denotes the replication padding operation with stride $S_{l}=P_{l}/2$ and $\mathrm{Unfold}_{P_{l},S_{l}}$ unfolds the padded sequence into patches of length $P_{l}$ using a stride of $S_{l}$.
+This operation produces a 3D patch tensor $\mathbf{X}_{u}^{l}\in\mathbb{R}^{C\times{N_{l}}\times{P_{l}}}$.
+Finally, EMPD projects $\mathbf{X}_{p}^{l}$ into a unified embedding space via a linear projection:
 $\mathbf{Z}^{l}=\mathrm{Projector}(\mathbf{X}_{p}^{l}),$ (5)
-This direct linear projection avoids redundant flattening operations, efficiently transforming patch-wise temporal features into a unified embedding space while preserving the hierarchical structure of multiscale patches.
+This direct linear projection efficiently maps patch-wise temporal features into a unified embedding space, avoiding redundant flattening operations while preserving the hierarchical structure of the multi-scale patches.
 Triad Interaction Block
-TIB is designed to model heterogeneous dependencies within the multiscale patch representations generated by EMPD, through joint capture of intra-patch, inter-patch, and cross-variable features.
-Given the input tensor $\mathbf{Z}^{l}\in{R}^{C\times{N_{l}}\times{D}}$, TIB integrates three complementary dependency types in a coherent framework.
-Specifically, TIB first processes intra-patch dependencies using depth-wise separable convolutions via $\mathbf{\zeta}_{intra}$ to capture fine-grained local patterns:
-$\mathbf{F}_{{intra}}^{l}=\mathbf{\zeta}_{intra}(\mathbf{Z}^{l}),$ (6)
-where $\mathbf{\zeta}_{intra}$ consists the depth-wise convolution and $\mathrm{Conv1d}$ to captures local temporal contexts and project features back to the embedding space.
-This operation preserves intra-patch continuity while extracting localized representations.
+The Triad Interaction Block(TIB) is designed to model heterogeneous dependencies within the multi-scale patch representations generated by EMPD, through joint capture of intra-patch, inter-patch, and cross-variable interactions.
+Given the input tensor $\mathbf{Z}^{l}\in\mathbb{R}^{C\times{N_{l}}\times{D}}$, TIB integrates these three complementary dependency types within a coherent framework.
+Specifically, TIB first processes intra-patch dependencies using depth-wise separable convolutions through operation $\mathbf{\zeta}_{\text{intra}}$ to capture fine-grained local patterns:
+$\mathbf{F}_{\text{intra}}^{l}=\mathbf{\zeta}_{\text{intra}}(\mathbf{Z}^{l}),$ (6)
+where $\mathbf{\zeta}_{\text{intra}}$ comprises depth-wise convolution and $\mathrm{Conv1d}$ to capture local temporal contexts and project features back to the embedding space.
+This operation preserves intra-patch continuity while effectively extracting local representations.
 For inter-patch dynamics $\mathbf{F}_{{intra}}^{l}$, TIB employs dilated convolutions with adaptive pooling via $\mathbf{\zeta}_{inter}$:
-$\mathbf{F}_{{inter}}^{l}=\mathbf{\zeta}_{inter}(\mathbf{F}_{{intra}}^{l}),$ (7)
-where $\mathbf{\zeta}_{inter}$ consists the dilated convolution and adaptive pooling to capture broader temporal contexts and patch-level information without increasing computational cost.
-Then TIB models dependencies between different features by generating adaptive weighting coefficients $\mathbf{G}^{l}\in\mathbf{R}^{{C}\times{1}}$ from globally aggregated inter-patch features $\bar{\mathbf{F}}_{{inter}}^{l}\in\mathbf{R}^{{1}\times{D}}$ as follows:
-$\mathbf{G}^{l}=\sigma\left(\mathrm{MLP}(\bar{\mathbf{F}}_{{inter}}^{l})\right),$
-$\mathbf{F}_{{cross}}^{l}=\mathbf{G}^{l}\odot\mathbf{F}_{{inter}}^{l},$ (8)
-where $\bar{\mathbf{F}}_{{inter}}^{l}$ denotes the global average of $\mathbf{F}_{{inter}}^{l}$, and the sigmoid function $\sigma$ act as a feature-specific gate.
-The weights $\mathbf{G}^{l}$ are then applied element-wise to $\mathbf{F}_{{inter}}^{l}$ through multiple operation $\odot$, which adaptively scales each feature’s contribution based on its own global relevance, enabling context-aware cross-variable interactions.
-Finally, three dependency representations are adaptively fused via learned gating weights, and the fused features are then integrated with residual connection and normalization:
-$\mathbf{F}^{l}_{{fused}}=\sum_{i=1}^{3}\mathbf{G}^{l}_{i}\odot\mathbf{F}^{l}_{i},$
-$\mathbf{F}^{l}_{{output}}=\mathrm{LayerNorm}(\mathbf{F}^{l}_{{fused}}+\mathbf{Z}^{l}),$ (9)
-where $\mathbf{G}^{l}_{i}$ is learned gating weights for each representations.
-This architecture enables TIB to dynamically balance contributions of different dependency types based on input characteristics, thus effectively capturing intricate dependencies, providing rich temporal features for multiscale prediction.
+$\mathbf{F}_{\text{inter}}^{l}=\mathbf{\zeta}_{\text{inter}}(\mathbf{F}_{\text{intra}}^{l}),$ (7)
+where $\mathbf{\zeta}_{\text{inter}}$ consists of the dilated convolution and adaptive pooling operations, capturing broader temporal contexts and patch-level information without increasing computational cost.
+TIB then models cross-variable dependencies by generating adaptive weighting coefficients $\mathbf{G}^{l}\in\mathbb{R}^{{C}\times{1}}$ as follows:
+$\mathbf{G}^{l}=\sigma\left(\mathrm{MLP}(\bar{\mathbf{F}}_{\text{inter}}^{l})\right),$
+$\mathbf{F}_{\text{cross}}^{l}=\mathbf{G}^{l}\odot\mathbf{F}_{\text{inter}}^{l},$ (8)
+where $\bar{\mathbf{F}}_{\text{inter}}^{l}$ denotes the global average of $\mathbf{F}_{\text{inter}}^{l}$, and the sigmoid function $\sigma$ acts as a feature-specific gate mechanism.
+$\mathbf{G}^{l}$ are applied element-wise to $\mathbf{F}_{\text{inter}}^{l}$ via the Hadamard product $\odot$, adaptively scaling each feature’s contribution based on its global relevance and enabling context-aware cross-variable interactions.
+Finally, the three dependency representations are adaptively fused using learned gating weights, and the resulting features are integrated with residual connection and normalization:
+$\mathbf{F}^{l}_{\text{fused}}=\sum_{i=1}^{3}\mathbf{G}^{l}_{i}\odot\mathbf{F}^{l}_{i},$
+$\mathbf{F}^{l}_{\text{output}}=\mathrm{LayerNorm}(\mathbf{F}^{l}_{\text{fused}}+\mathbf{Z}^{l}),$ (9)
+where $\mathbf{G}^{l}_{i}$ represents the learned gating weights for the $i$-th representations.
+This architecture enables TIB to dynamically balance the contributions of different dependency types based on the temporal characteristics of input data.
+Consequently, it effectively captures intricate dependencies and provides rich temporal features for multi-scale prediction.
 Models DMSC(Ours) TimeMixer iTransformer PatchTST Dlinear TimesNet Autoformer TimeXer PatchMLP TimeKAN AMD
@@
 Adaptive Scale Routing Mixture-of-Experts
-To address the limitations of static fusion mechanisms in multiscale forecasting, we propose the Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) block.
-This dynamic prediction head models temporal dependencies at different horizons via hierarchical experts (global-shared and local-specialized experts) and adaptively weights scale-specific predictions based on temporal patterns.
-Specifically, ASR-MoE comprises two expert groups to handle distinct temporal granularities.
-Global experts $\mathcal{E}^{g}=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{m}\}$ capture common long-term dependencies through deep nonlinear transformations, modeling persistent trends and long-periodic patterns that commonly existed in time series.
-Local experts $\mathcal{E}^{l}=\{\mathcal{L}_{1},\mathcal{L}_{2},...,\mathcal{L}_{n}\}$ specialize in different short-term variations through shallower networks, detecting short-periodic patterns and high-frequency fluctuations that dynamically varied across stages.
-This explicit decoupling enables specialized handling of complex dynamics while maintaining higher parameter efficiency.
-For weights calculation of each scale’s experts, a dynamic routing mechanism is proposed to assign input-dependent weights to experts, balancing global and local dependencies.
-Given a scale-specific feature $\mathbf{F}_{l}$, the router generates weights as follows:
+To address the limitations of static fusion mechanisms in multi-scale forecasting, we propose the Adaptive Scale Routing Mixture-of-Experts (ASR-MoE) block.
+This dynamic prediction head captures temporal dependencies at different horizons through a hierarchy of experts (global-shared and local-specialized experts) and adaptively aggregates scale-specific predictions based on temporal patterns.
+Specifically, ASR-MoE incorporates two distinct expert groups to handle different temporal granularities.
+Global experts $\mathcal{E}^{g}=\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{m}\}$ capture common long-term dependencies through deeper networks, modeling persistent trends and long-periodic patterns that commonly existed in time series.
+Local experts $\mathcal{E}^{l}=\{\mathcal{L}_{1},\mathcal{L}_{2},...,\mathcal{L}_{n}\}$ capture diverse short-term variations via shallower networks, detecting short-periodic patterns and high-frequency fluctuations that vary dynamically across stages.
+This explicit decoupling enables specialized handling of complex temporal dynamics while maintaining high parameter efficiency.
+To compute the weighting for experts at each scale, a dynamic routing mechanism assigns input-dependent weights, thereby balancing the contributions of global and local dependencies.
+Given a scale-specific feature $\mathbf{F}_{l}$, the router generates weights, and only the top-$K$ local experts are activated via sparse routing:
+$\mathbf{\hat{\Omega}}^{L}_{l},\mathcal{I}=\begin{cases}\mathbf{\Omega}^{L}_{l},&\mathbf{\Omega}^{L}_{l}\in\mathrm{Top-K}(\mathbf{\Omega}_{l},K)\\ 0,&\text{otherwise}\end{cases}$
+$\mathbf{\Omega}_{G}=\mathrm{Sigmoid}(\mathrm{MLP}(\mathbf{F}_{l})),$
 $\mathbf{\Omega}_{l}=\mathrm{Softmax}(\mathrm{MLP}(\mathbf{F}_{l})),$ (10)
-where the weight matrix $\mathbf{\Omega}$ are divided into global weights $\mathbf{\Omega}^{G}$ and local weights $\mathbf{\Omega}^{L}$.
-For computational efficiency, we activate only the top-$K$ local experts via sparse routing:
-$\mathbf{\hat{\Omega}}^{L}_{l},\mathcal{I}=\mathrm{Top-K}(\mathbf{\Omega}^{L}_{l},K),$ (11)
-where $\mathcal{I}$ denotes indices of selected experts.
-Finally, a temporal-aware weighting module fuses outputs from all scales by historical scale importance. Scale-specific predictions $\hat{\mathbf{Y}}^{l}$ are fused using time-dependent weights:
-$\mathbf{w}=\mathrm{Softmax}(\mathrm{MLP}(\bigoplus^{L}_{l=1}\phi(\mathbf{F}_{l})\cdot\mathbf{w}_{hist})),$ (12)
-where $\phi(\cdot)$ denotes temporal descriptors, $\mathbf{w}_{hist}$ denotes historical weighting memory, and $\bigoplus$ means concatenation. The final prediction integrates multi-scale outputs:
+where $\mathbf{\Omega}_{g}$ and $\mathbf{\Omega}_{l}$ represent the global and local expert weight matrices, respectively.
+$\mathcal{I}$ denotes the set of indices for the selected experts.
+Finally, a temporal-aware weighting module fuses outputs from all scales by historical scale importance.
+The scale-specific predictions $\hat{\mathbf{Y}}^{l}$ are fused using time-dependent weights:
+$\mathbf{w}=\mathrm{Softmax}\left(\mathrm{MLP}\left(\bigoplus^{L}_{l=1}\phi(\mathbf{F}_{l})\cdot\mathbf{w}_{\text{hist}}\right)\right),$ (11)
+where $\phi(\cdot)$ denotes temporal descriptors, $\mathbf{w}_{\text{hist}}$ represents the historical weighting memory, and $\bigoplus$ denotes the concatenation operation.
+The final prediction is obtained by integrating the multi-scale outputs:
 $\hat{\mathbf{Y}}=\sum_{l=1}^{L}\mathbf{w}_{l}\left(\sum_{M}^{m=1}{\mathbf{\Omega}^{G}_{m}\mathcal{G}_{m}}(\mathbf{F}_{l})+\sum_{n\in{\mathcal{I}}}{\hat{\mathbf{\Omega}}^{L}_{n}\mathcal{L}_{n}}(\mathbf{F}_{l})\right),$ (12)
-Meanwhile, to ensure balanced expert utilization, ASR-MoE incorporates a balance loss:
+To ensure balanced expert utilization, ASR-MoE incorporates an auxiliary balance loss:
 $\mathcal{L}_{{balance}}=-\lambda\mathbf{E}[\sum\mathbf{\Omega}_{j}log\mathbf{\Omega}_{j}],$ (13)
-and the overall loss function of DMSC is as follows:
+The overall loss function for DMSC is defined as follows:
 $\mathcal{L}=\mathcal{L}_{pred}+\mathcal{L}_{{balance}},$ (14)
 where $\mathcal{L}_{pred}$ denotes the Mean Squared Error (MSE) loss.
-By combining hierarchical expert specialization, dynamic routing, and temporal-aware weighting, ASR-MoE enables the proposed framework to adaptively prioritize relevant scales and experts, achieving a synergy between long-term trend projection and short-term detail refinement in TSF.
+By integrating hierarchical expert specialization, dynamic routing, and temporal-aware weighting, ASR-MoE adaptively prioritizes relevant scales and experts, achieving a synergistic balance between long-term trend capture and short-term detail refinement in time series forecasting.
 Experiments
-To comprehensively verify the performance and effectiveness of DMSC, we conduct extensive experiments on 13 real-world TSF benchmarks spanning multiple domains and temporal resolutions.
+To comprehensively evaluate the performance and effectiveness of the proposed DMSC framework, we conduct extensive experiments on 13 real-world TSF benchmarks across multiple domains and temporal resolutions.
 Datasets.
 For long-term forecasting, we conduct experiments on nine well-established benchmarks, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, ECL, Traffic, Weather and Solar.
-For short-term forecasting, we experiment on PEMS(PEMS03, PEMS04, PEMS07, PEMS08).
+For short-term forecasting, we utilize the PEMS(PEMS03, PEMS04, PEMS07, PEMS08) dataset.
 Baselines.
@@
 Main Results
-The main results are listed in Table 1 for long-term forecasting and Table 2 for short-term forecasting, in which the lower MSE and MAE indicates better forecasting performance.
-Our proposed DMSC framework demonstrates consistent SOTA performance across all 13 benchmarks, achieving the lowest MAE and MSE in most experimental configurations.
-These results confirm robust effectiveness and generality of DMSC for both long-term and short-term forecasting tasks.
-Notably, multi-scale architectures like TimesNet and PatchMLP exhibit instability due to their inflexible fusion mechanisms that fails to fully leverage captured multiscale features.
-And the preferable performance of TimeXer and iTransformer underscores the effect of the patch-wise and variate-wise strategy in information representation, but their fixed decomposition schemes inherently constrain the model’s potential.
-Compared to iTransformer and Dlinear, DMSC has a considerable improvement in data with more variate(Traffic, PEMS), demonstrating the necessity of jointly dependency modeling.
-Collectively, the above observations validate DMSC’s efficacy in addressing the core challenges of time series forecasting.
+The main results for long-term and short-term forecasting are presented in Table 1 and Table 2, in which the lower MSE and MAE values indicates superior forecasting performance.
+The proposed DMSC framework demonstrates consistent SOTA performance across all 13 benchmarks, achieving the lowest MAE and MSE in most experimental settings.
+These results confirm the robust effectiveness and generalizability of DMSC for both long-term and short-term forecasting tasks.
+Notably, multi-scale architectures like TimesNet and PatchMLP exhibit performance instability, which we attribute to their inflexible fusion mechanisms that fails to fully leverage the captured multi-scale features.
+The competitive performance of TimeXer and iTransformer underscores the effectiveness of patch-wise and variate-wise strategy for information representation, but their fixed decomposition schemes inherently constrain the model’s potential.
+Furthermore, compared to iTransformer and Dlinear, DMSC achieves considerable improvement in data with more variate(Traffic, PEMS), highlighting the necessity of jointly dependency modeling.
+Collectively, the above observations validate the efficacy of DMSC in addressing the core challenges of time series forecasting.
 Model Analysis
@@
 Table 6: Ablation study on ASR-MoE to assess the importance of expert specialization (global/local) and dynamic fusion strategies for prediction results.
-For EMPD, employing predefined static patches or raw data embedding instead of input-adaptive decomposition degraded performance.
-This demonstrates the necessity of patch-wise decomposition and lightweight neural controllers for granularity adjustment.
-For TIB, replacing joint modeling with standard convolution layers or only intra-patch features resulted in performance drops, validating the need for comprehensive and joint modeling of triadic dependencies.
-Meanwhile, ablation of dynamic fusion mechanism in TIB shows its effectiveness to dynamically fuse these heterogeneous dependencies, as static handling of such relationships undermines holistic dependency capture across different scales.
-For ASR-MoE, substituting our adaptive prediction mechanism with either a single prediction head or summation of multiple linear heads yields substantial performance declines, which demonstrates that single heads inadequately utilize multiscale features, and simple aggregation of multiple heads overlooks the differential importance of scale-specific dependencies.
-Moreover, removing either global or local experts leads to performance degradation, which demonstrates the necessity of leveraging distinct experts for different temporal patterns.
-The elaborate design in ASR-MoE for collective utilization of specialized experts can effectively harness diverse temporal dynamics.
-Collectively, these ablation studies confirm that dynamic and multiscale modeling is central to DMSC.
-All of three blocks and the proposed progressive cascade architecture together enable robust multiscale modeling and hierarchical feature learning.
-The consistent performance drops from ablating any component validate the integrity of DMSC’s architecture.
+For EMPD, replacing the input-adaptive decomposition with either predefined static patches or raw data embedding leads to performance degradation.
+This demonstrates the necessity of adaptive patch-wise decomposition and its lightweight network for dynamic granularity adjustment.
+For TIB, replacing joint modeling with standard convolution layers or only intra-patch features resulted in performance drops, validating the necessity of comprehensive joint modeling of triadic dependencies.
+Similarly, ablation of dynamic fusion mechanism in TIB confirms its effectiveness in fusing heterogeneous dependencies, as static handling undermines holistic cross-scale dependency capture.
+For ASR-MoE, replacing the adaptive prediction mechanism with a single prediction head or summed multiple linear projections leads to substantial performance declines.
+This indicates that single heads inadequately utilize multi-scale features, while simplistic aggregation overlooks scale-specific dependency importance.
+Moreover, removing either global or local experts degrades performance, demonstrating the necessity of maintaining distinct experts for different temporal patterns.
+The elaborate design of ASR-MoE effectively harnesses diverse temporal dynamics through collective specialized expert utilization.
+Collectively, these ablation studies confirm that dynamic and multi-scale modeling is central to DMSC.
+All of three blocks and the proposed progressive cascade architecture together enable robust multiscale modeling and hierarchical feature learning, with consistent performance drops from ablating any component validate architectural integrity of DMSC.
 Figure 2: Visualization of different experts on ETTh1 datasets, the look-back and prediction length are set to 96.
 ASR-MoE Expert Specialization.
-To verify the specialization of ASR-MoE experts, we conducted activation pattern visualization across different experts.
+To verify the specialization of ASR-MoE experts, we visualize their activation patterns across different temporal experts.
 As shown in Fig.
-2, distinct specialization profiles emerge: global experts exhibit pronounced sensitivity to persistent trends and long-periodic patterns, while local experts demonstrate acute responsiveness to short-periodic patterns and high-frequency fluctuations.
-This specialization demonstrates adaptively knowledge learning for different experts.
+2, distinct specialization profiles emerge: global experts exhibit heightened sensitivity to persistent trends and long-periodic patterns, whereas local experts demonstrate acute responsiveness to short-periodic patterns and high-frequency fluctuations.
+This observed specialization demonstrates the adaptive acquisition of specialized knowledge by different experts.
 Increasing Look-Back Length.
-Theoretically, longer input sequences provide richer historical information, potentially enhancing the model’s ability to capture long-term trends and complex multiscal dependencies.
-Our DMSC framework is specifically designed to leverage this, with its dynamic multiscale coordination capability inherently enabling adaptive modeling and effective utilization of information across varying look-back lengths.
-So we evaluate DMSC with look-back length from {48, 96, 192, 336, 720} on Electricity dataset.
+Theoretically, longer input sequences provide richer historical information, which can enhance the model’s ability to capture long-term trends and complex multi-scale dependencies.
+Our DMSC framework is specifically designed to leverage this characteristic, as its dynamic multi-scale coordination capability inherently facilitates adaptive modeling and effective information utilization across varying look-back lengths.
+So we evaluate DMSC with look-back lengths selected from {48, 96, 192, 336, 720} on Electricity dataset.
 As shown in Fig.
-3, DMSC consistently improves performance with longer historical sequences while outperforms other baselines, demonstrating its capacity to extract richer temporal representations from extended contexts and effectiveness in integrating multiscale temporal patterns.
+3, DMSC consistently achieves improved performance with longer historical sequences while outperforming other baseline methods.
+This demonstrates its capacity to extract richer temporal representations from extended contexts and effectively integrate multi-scale temporal patterns.
 Figure 3: Forecasting results with varying look-back length on Electricity dataset. Look-back lengths are set to {48, 96, 192, 336, 720}, and prediction length is fixed to 96.
 Efficiency Analysis
-We rigorously evaluate the memory usage and training time of DMSC against other SOTA baselines on Weather(21 variates) and Traffic datasets(862 variates).
+We rigorously evaluate the memory usage and training time of DMSC against other SOTA baselines on Weather(21 variates) and Traffic(862 variates) datasets.
 As demonstrated in Fig.
@@
 5, DMSC consistently achieves superior efficiency in both memory usage and training time while maintaining competitive forecasting accuracy.
-Crucially, DMSC demonstrates near-linear scalability to long sequences—significantly outperforming Transformer-based models that exhibit quadratic complexity and MLP-based approaches with parameter explosion.
-These efficiency gains are particularly pronounced when handling high-dimensional multivariate data (Traffic), where DMSC’s cross-variable interaction avoids the computational overhead of exhaustive pairwise attention while maintaining modeling capacity.
-Our framework thus establishes a balance of accuracy, latency, and memory usage for practical deployment scenarios.
+Crucially, DMSC demonstrates near-linear scalability to long sequences, as outperforming Transformer-based models with quadratic complexity and MLP-based approaches susceptible to parameter explosion.
+These efficiency gains are particularly pronounced when processing high-dimensional multivariate data (Traffic), where the cross-variable interaction in TIB avoids the computational overhead of exhaustive pairwise attention while preserving modeling capacity.
+DMSC thus establishes a balance of accuracy, latency, and memory usage for practical deployment scenarios.
 Figure 4: Model efficiency analysis under 96-look-back length and 96-prediction length on Weather(21 variates) datasets. Batch size is set to 128.
@@
 Conclusion
-This paper introduces the Dynamic Multi-Scale Interaction (DMSC) framework, a novel framework that can advance TSF through dynamic multiscale coordination.
-DMSC has three key contributions: 1) an Embedded Multi-Scale Patch Decomposition(EMPD) is introduced to dynamically decompose time series into hierarchical patches to eliminate predefined and fixed scale constraints; 2) Triad Interaction Block (TIB) is proposed to jointly model intra-patch, inter-patch, and cross-variable dependencies, thus forming a coarse-to-fine feature pyramid through progressive cascade layers; 3) an Adaptive Scale Routing MoE (ASR-MoE) is proposed to dynamically fuse multiscale predictions via temporal-aware weighting of global and local experts.
-Extensive experiments on thirteen benchmarks demonstrate that DMSC can achieve SOTA performance and superior efficiency.
-Future work will extend the framework to multi-tasks learning and explore optimizations on intricate dataset environments, improving its multi-scenario applicability for real world.
+This paper presents the Dynamic Multi-Scale Coordination (DMSC) framework, a novel approach that advances time series forecasting through comprehensive dynamic multi-scale coordination.
+DMSC has three key contributions: 1) Embedded Multi-Scale Patch Decomposition(EMPD) dynamically decomposes time series into hierarchical patches, eliminating predefined scale constraints through input-adaptive granularity adjustment; 2) Triad Interaction Block (TIB) jointly models intra-patch, inter-patch, and cross-variable dependencies, forming a coarse-to-fine feature pyramid through progressive cascade layers; 3) Adaptive Scale Routing MoE (ASR-MoE) dynamically fuses multi-scale predictions via temporal-aware weighting of specialized global and local experts.
+Extensive experiments on thirteen benchmarks demonstrate that DMSC achieves SOTA performance and superior efficiency.
+Future work will extend the framework to multi-task learning and explore optimizations on complex real-world environments, thereby enhancing its applicability across diverse scenarios.
 Appendix A Implementation Details
@@
 The embedding dimension determines the richness of feature representations.
-We set $d_{\mathrm{model}}\in\{64,128,256,512\}$ to evaluate model’s performance on ECL and Weather dataset.
-The results are shown in Fig.
+We evaluate the model’s performance on ECL and Weather datasets with embedding dimensions $d_{\mathrm{model}}\in\{64,128,$ 256,512}256,512\}.
+The results in Fig.
 6(a) and Fig.
-6(b), we choose the optimal dimension at 128 / 256 - smaller embedding impairs multiscale feature separation, while larger configuration yield marginal gains with higher memory cost, indicating dimensionality saturating.
+6(b) show that the optimal dimension is at 128 / 256, as smaller embedding impairs multi-scale feature separation, while larger configurations yield marginal gains with higher memory cost, indicating dimensionality saturating.
 (a) Performance on Electricity dataset.
@@
 EMPD Patch Length Decay.
-EMPD employs exponentially decaying patch lengths to capture multiscale dependencies.
-We vary the decay rate (2, 3, 4) and find that a decay factor of 2 yields the best results from Fig.
-7(a) on weather dataset.
-It can be seen that a decay factor of 2 yields the best forecasting performance, which balances the preservation of sufficient local details in fine-grained layers against the extraction of broader trends in coarse-grained layers.
+EMPD employs exponentially decaying patch lengths across layers to capture multi-scale dependencies.
+We evaluate the impact of the EMPD decay rate by testing values of $\tau\in\{2,3,4\}$ on Weather dataset.
+As shown in Fig.
+7(a), a decay factor of $\tau=2$ achieves the best forecasting performance, which balances the preservation of sufficient local details in fine-grained layers with the extraction of broader trends in coarse-grained layers.
 Number of Progressive Cascade Layers.
-The number of cascade layers ($l\in\{1-5\}$) affects the hierarchical feature extraction.
+The number of progressive cascade layers ($l\in\{1,2,3,4,5\}$) influences hierarchical feature extraction.
 As shown in Fig.
-7(b), three layers can achieve peak efficiency - shallower stacks (1 - 2 layers) preserve perfect accuracy via residual propagation but miss fine-grained interactions, while deeper configurations (4 - 5 layers) introduce diminishing returns with much more latency but fewer improvement.
+7(b), three layers can achieve optimal efficiency, shallower stacks (1 - 2 layers) fail to capture fine-grained interactions, while deeper configurations (4 - 5 layers) introduce substantially increased latency with diminishing performance returns.
 Appendix C Full Results